### v1

# A new approach to principal–agent problems with volatility control

Abstract
The recent work by Cvitanić, Possamaï, and Touzi (2018) [9] presents a general approach for continuous-time principal–agent problems, through dynamic programming and second-order backward stochastic differential equations (BSDEs). In this paper, we provide an alternative formulation of the principal–agent problem, which can be solved simply by relying on the theory of BSDEs. This reformulation is strongly inspired by an important remark in [9], namely that if the principal observes the output process $X$ in continuous-time, she can compute its quadratic variation pathwise. While in [9], this information is used in the contract, our reformulation consists in assuming that the principal could directly control this process, in a ‘first-best’ fashion. The resolution approach for this alternative problem actually follows the line of the so-called ‘Sannikov’s trick’ in the literature on continuous-time principal–agent problems, as originally introduced by Sannikov (2008) [29]. We then show that the solution to this ‘first-best’ formulation is identical to the solution of the original problem. More precisely, using the contract form introduced in [9] as penalisation contracts, we highlight that this ‘first-best’ scenario can be achieved even if the principal cannot directly control the quadratic variation. Nevertheless, we do not have to rely on the theory of 2BSDEs to prove that such contracts are optimal, as their optimality is ensured by showing that the ‘first-best’ scenario is achieved. We believe that this more straightforward approach to solve continuous-time principal–agent problems with volatility control will facilitate the dissemination of these problems across many fields, and its extension to even more intricate problems.
Keywords. Principal–agent problems, volatility control, moral hazard, first-best, BSDEs.
AMS 2020 subject classifications. Primary: 91B43; secondary: 91B41, 93E20.
1 Introduction
Contract theory, and more precisely principal–agent problems, offer a relevant framework for the study of optimal incentives between agents, especially with information asymmetry. In standard principal–agent problems, a principal (she) is imperfectly informed about the actions of an agent (he). To incentivise the agent to act in her best interest, she can offer him a contract, usually represented by a terminal payment $\xi$ at a fixed finite time horizon $T>0$. The main mathematical challenge of these problems is to determine an optimal contract, in the sense that it maximises the utility of the principal, while that of the agent is held at least to a given level. The design of such a contract raises first of all an issue of information: the contract can only depend on what is known, or observable, by the principal. We focus in this paper on the moral hazard case, i.e. when the effort of the agent is hidden and therefore not contractible.
For a long time, these problems were only considered in discrete-time or static settings, in particular with the works of Mirrlees in [23] (1976) and [24] (1999), or the seminal papers of Holmström (1979) [15] and Grossman and Hart (1983) [12]. However, discrete-time models are actually challenging to solve, especially because the existence of an optimal contract in these settings is more the exception than the rule, even in very simple and natural examples. The theory then regained momentum with the first continuous-time formulation of the problem by Holmström and Milgrom (1987) [16], and later by Sannikov (2008) [29], using tools from stochastic control theory. Through these works, it is established that, if the agent only controls the drift of a stochastic process $X$, continuously observable by the principal, then the optimal contract $\xi$ should have the following form,
$\xi=y_{0}-\int_{0}^{T}{\cal H}(t,X_{t},Z_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\cdot\mathrm{d}X_{t},$ (1.1)
where ${\cal H}$ here is the (simplified) agent’s Hamiltonian for drift control. In the previous contract form, the constant $y_{0}\in\mathbb{R}$ and the process $Z$ have to be optimally chosen by the principal, in order to maximise her utility and ensure the participation of the agent in the contractual agreement. More precisely, the choice of $y_{0}\in\mathbb{R}$ is usually dictated by the agent’s participation constraint, while the process $Z$ will incentivise the agent to perform appropriate effort on the drift of the output process. This form of contract naturally comes from the study of the agent’s continuation utility, which, in the case of drift control only, can be related to the solution of a backward stochastic differential equation (BSDE for short). This formal link between the agent’s continuation utility and the contract itself was introduced by Sannikov (2008) [29], and thus usually referred to as ‘Sannikov’s trick’ in the literature on continuous-time principal–agent problems, but has recently been revisited in a more theoretical way by Possamaï and Touzi (2020) [27]. These problems have subsequently been extended in many directions, for example to consider multi-agent and/or multi-principal problems, a continuum of agents with mean-field interactions, more general output process $X$, more intricate information asymmetry between the two parties, to only cite a few.
In terms of theoretical developments, the most significant breakthrough is thanks to Cvitanić, Possamaï, and Touzi (2018) [9], who have developed a general and rigorous theory that allows to address a wide spectrum of principal–agent problems, especially when the agent can also control the volatility of the output process. As explained in [9], moral hazard problems in continuous-time in which the agent controls the volatility of the output are notoriously harder to study. Indeed, when the agent is allowed to also control the volatility of the output process, his continuation utility cannot be related to the solution of a BSDE anymore, one has to rely now on the theory of backward stochastic differential equations of the second order (2BSDE for short), introduced by Soner, Touzi, and Zhang (2012) [30] and further developed by Possamaï, Tan, and Zhou (2018) [28]. One important observation made in [9] which will also be at the heart of our approach is that, if the principal has access to the trajectory of the output process $X$, then she can deduce its quadratic variation process $\langle X\rangle$ in a pathwise manner, as demonstrated by Bichteler (1981) [4]. This motivates the authors to introduce a new form of contracts, namely
$\xi=y_{0}-\int_{0}^{T}{\cal H}(t,X_{t},Z_{t},\Gamma_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\cdot\mathrm{d}X_{t}+\dfrac{1}{2}\int_{0}^{T}{\rm Tr}\big{[}\Gamma_{t}\mathrm{d}\langle X\rangle_{t}\big{]}.$ (1.2)
Such contracts are indexed on $X$ as before, but also on its quadratic variation $\langle X\rangle$ through an additional parameter $\Gamma$, which should incentivise the agent to perform appropriate efforts on the volatility. Note that the agent’s Hamiltonian ${\cal H}$ here should be different from the one introduced in the previous contract form, precisely to take into account volatility control. When restricting the study to this form of contracts, it is straightforward to show, mostly using Ito’s formula, that the agent’s optimal effort will naturally be given by the maximiser of the Hamiltonian. One can then proceed by solving the principal’s optimal control problem, namely by writing the Hamilton-Jacobi-Bellman (HJB for short) equation corresponding to her stochastic control problem, which becomes relatively standard under the restriction to the previous form of contracts. These first two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the third and last step requires to prove that the restriction to such form of contracts is without loss of generality. The approach provided in [9] to prove this main result relies on the recent and sophisticated theory of 2BSDEs. The aim of this paper is precisely to present an alternative approach to solving this type of problems, which avoids relying on this theory. For the sake of comparison, we will study the same general model, whose original formulation will be recalled in Section 2.
Despite the challenges raised when studying principal–agent problems with volatility control, especially as the optimality of revealing contracts relies on deep results from the recent theory of 2BSDEs, it has subsequently been extended and applied in many different situations. We can mention in a non-exhaustive way the applications related to finance by the same authors in [8], by Cvitanić and Xing (2018) [7] or more recently by Baldacci and Possamaï (2022) [3]; the works by Aïd, Possamaï, and Touzi (2022) [1], Élie, Hubert, Mastrolia, and Possamaï (2021) [11] and Aïd, Kemper, and Touzi (2023) [2] for applications related to the energy sector; as well as other various extensions, e.g., the works of Mastrolia and Possamaï (2018) [22], Hernández Santibáñez and Mastrolia (2019) [14], Hu, Ren, and Touzi (2019) [17], Hubert (2023) [18] or Keppo, Touzi, and Zuo (2024) [21]. Naturally, the study of applications fitting into the general framework proposed in [9] can be done in a fairly straightforward way, directly using the results, as done in [8, 1]. Nevertheless, the recent research mentioned above has also attempted to generalise principal–agent problems, especially to consider many agents (see [18, 2]), or even a continuum (see [11]). If one follows the approach developed in [9], these extensions require to introduce the so-called multidimensional or mean-field 2BSDEs, and study, or assume, the well-posedness of such systems. However, while the theory on (one-dimensional) 2BSDEs is sufficiently developed to address relatively general principal–agent problems, there is currently no definition nor existence results on multidimensional/mean-field 2BSDEs. This further motivates us to present, in this paper, an alternative approach to solving the principal–agent problem studied in [9]. We believe that this novel approach, which avoids relying on the theory of 2BSDEs, will greatly facilitate the study of relevant generalised principal–agent problems, especially with many agents, or non-continuous output processes.
Before introducing the approach developed in this paper, we wish to provide more intuition on the problem under consideration via a short detour through the so-called first-best case. When studying principal–agent problems with moral hazard, it is quite common to compare the results with this benchmark scenario, in which the principal can directly optimise the agent’s effort. This problem is usually easier to solve, as one only has to consider the principal’s stochastic control problem under the agent’s participation constraint. It may happen in the moral-hazard case, sometimes also referred to as second-best, that the principal can actually deduce, through her observation, the effective efforts of the agent. Under this assumption, the two problems are then equivalent if the principal can design a contract which strongly penalises the agent whenever he deviates from the optimal effort recommended by the principal. Typically, the principal will offer to the agent a minimal [^1] Minimal in the sense that it would just be sufficient to ensure the participation of agent. compensation if he follows the recommendation, and threaten him with a high penalty if he deviates from the optimal effort. This can usually be done if no strong assumptions are made on the contract, but can fail to be implemented in cases for example with limited liability, in which the contract is required to be positive.
As highlighted before, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] relies in particular on the observation that, if the principal have access to the trajectory of the output process $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise, and thus design contracts of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation of the original principal–agent problem with volatility control is that, since the principal can deduce the quadratic variation $\langle X\rangle$, she could probably force the agent to choose his efforts in order to achieve an optimal quadratic variation. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original problem, in which the principal enforces a contract $\xi$ as well as a quadratic variation process. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal. This ‘first-best’ alternative formulation of the problem is described with more details in Section 3.1. When carefully studying the agent’s problem in this reformulation of the original problem, one can remark that, since the quadratic variation is a priori fixed by the principal, the agent’s weak formulation boils down to choosing equivalent probability measures by means of Girsanov theorem. For a fixed quadratic variation process and any contract $\xi$, the agent’s continuation utility can therefore be represented by means of a BSDE, instead of a 2BSDE in the original formulation. More informally, in this reformulated problem, one can apply the so-called ‘Sannikov’s trick’ to obtain a relevant representation of the contract $\xi$, which is then similar to the optimal form of contracts in principal–agent problems with drift control only, namely (1.1). Finally, the principal’s problem boils down to a more standard stochastic control problem, in the sense that she should maximise her expected utility by optimally choosing the contract’s parameters, namely a constant $y_{0}\in\mathbb{R}$ and an appropriate process $Z$, as well as the quadratic variation process.
We wish to insist on the fact that the aforementioned approach to solve the ‘first-best’ alternative formulation of the original problem is quite straightforward, as it follows the ‘Sannikov’s trick’. More importantly, we establish that the principal’s value in this reformulated problem coincides with the one obtained in [9]. More precisely, from the principal’s point of view, optimising over $\Gamma$ in the contract form (1.2) or directly optimising over admissible quadratic variation lead to the exact same results. In particular, the agent’s optimal effort as well as his utility remain unchanged in both formulations. This result is actually quite intuitive: as contracts of the form (1.2) introduced in [9] are optimal when the principal observes the quadratic variation, they should allow her to achieve the expected utility she would have if she could directly control the quadratic variation process. In particular, this result confirms that contracts of the form (1.2) are also optimal when the agent’s efforts on the volatility can actually be determined by the principal through her observation of the quadratic variation. This equivalence between the two formulations is detailed in Section 3.3, using penalisation contracts.
As already mentioned before, the moral-hazard case is equivalent to its first-best counterpart if the principal can deduce the agent’s control from her observations and design a penalisation contract. Indeed, even if the principal observes the agent’s control, she may not have enough bargaining power to actually force the agent to perform the optimal controls. Showing existence of penalisation contracts thus ensures that the two frameworks are equivalent. Similarly, in our problem, even if the principal observes the quadratic variation of the output process, she may not be able to actually force the agent to follow the recommended quadratic variation. We therefore need to show the existence of such penalisation contracts, which allow the principal to achieve her ‘first-best’ utility even if she cannot directly control the quadratic variation. It is quite clear from the previous reasoning that the contract form (1.2) introduced in [9] will perfectly serve this purpose. We nevertheless wish to emphasise that, although we will use the contract form (1.2) as penalisation contracts, we never, with our approach, use the 2BSDE theory to prove optimality of such contracts, as it naturally comes from the fact that these contracts achieve the best possible utility for the principal. In other words, instead of proving via the theory of 2BSDE that considering contracts of the form (1.2) is without loss of generality from the principal’s point of view, one can simply show that these contracts allow her to reach her ‘first-best’ value.
To summarise, this paper provides an alternative approach to solve principal–agent problems with volatility control. More precisely, we first solve the ‘first-best’ formulation of the original problem by means of the theory of BSDEs, and then show that the contract form studied in [9] allows to achieve this first-best scenario, and is therefore naturally optimal. In other words, this alternative approach allows to prove that the contract form introduced in [9] is optimal, relying only on the theory of BSDEs. Although this approach ‘only’ achieves the same results as those presented by Cvitanić, Possamaï, and Touzi (2018) [9], we believe that this alternative, simpler approach will enable better dissemination of continuous-time principal–agent problems with volatility control in various fields across mathematical finance, economics, management sciences, operation research; fostering their use in many relevant applications. Furthermore, as this new approach only relies on the theory of BSDEs, we believe that it will facilitate the study of relevant extensions of principal–agent problems, for example when considering such a Stackelberg game with many agents, as well as its mean-field limit, when the principal can contract with a continuum of agents. This more straightforward approach could also simplify the extension of principal–agent problems to more general non-continuous output processes.
The structure of the paper is as follows. We first introduce in Section 2 the continuous-time principal–agent problem considered by Cvitanić, Possamaï, and Touzi (2018) [9]. We also briefly describe the main steps of the approach they developed to solve the problem, highlighting the main challenges posed by this approach. Section 3 details the main results of this paper: we first introduce in Section 3.1 the alternative ‘first-best’ reformulation of the original principal–agent problem, we then characterise in Section 3.2 the optimal form of contracts, relying only on the theory of BSDEs, and we finally prove in Section 3.3 that the solution of the reformulated problem coincides with the solution of the original problem. To further highlight the link between the two formulations, we provide in Section 4 some illustrative examples in which the solution can be explicitly characterised. Finally, Section 5 concludes by providing further related research directions.
Notations.
Let :=$\mathbb{N}^{\star}:=\mathbb{N}\setminus\{0\}$ be the set of positive integers, and recall that $T>0$ denotes some maturity fixed in the contract. For any $(\ell,c)\in\mathbb{N}^{\star}\times\mathbb{N}^{\star}$, $\mathbb{M}^{\ell,c}$ will denote the space of ×$\ell\times c$ matrices with real entries. When $\ell=c$, we let $\mathbb{M}^{\ell}:=\mathbb{M}^{\ell,\ell}$, and denote by $\mathbb{S}_{+}^{\ell}$ the set of $\ell\times\ell$ non-negative definite (or positive semi-definite) symmetric matrices with real entries. For any $d\in\mathbb{N}^{\star}$, let ${\cal C}([0,T],\mathbb{R}^{d})$ denote the set of continuous functions from $[0,T]$ to $\mathbb{R}^{d}$. The set ${\cal C}([0,T],\mathbb{R}^{d})$ will always be endowed with the topology associated to the uniform convergence on the compact $[0,T]$. For a probability space of the form $\Omega:={\cal C}([0,T],\mathbb{R}^{d})$ and an associated filtration $\mathbb{F}$, we will have to consider processes $\psi:[0,T]\times{\cal C}([0,T],\mathbb{R}^{k})\longrightarrow E$, taking values in some Polish space $E$, which are $\mathbb{F}$-optional, i.e. ${\cal O}(\mathbb{F})$-measurable where ${\cal O}(\mathbb{F})$ is the so-called optional $\sigma$-field generated by $\mathbb{F}$-adapted right-continuous processes. In particular, such a process $\psi$ is non-anticipative in the sense that $\psi(t,x)=\psi(t,x_{\cdot\wedge t})$, for all $t\in[0,T]$ and $x\in{\cal C}([0,T],\mathbb{R}^{k})$. Furthermore, for any filtration $\mathbb{F}:=({\cal F}_{t})_{t\in[0,T]}$ and probability measure $\mathbb{P}$, we will denote by:
  • $\bullet$
    $\mathbb{F}^{\mathbb{P}}:=({\cal F}_{t}^{\mathbb{P}})_{t\in[0,T]}$ the usual $\mathbb{P}$-completion of $\mathbb{F}$;
  • $\bullet$
    $\mathbb{F}^{+}:=({\cal F}_{t}^{+})_{t\in[0,T]}$ the right limit of $\mathbb{F}$;
  • $\bullet$
    $\mathbb{F}^{\mathbb{P}^{+}}:=({\cal F}_{t}^{\mathbb{P}^{+}})_{t\in[0,T]}$ the augmentation of $\mathbb{F}$ under $\mathbb{P}$.
Finally, for technical reasons, we work under the classical ZFC set-theoretic axiom. This is a standard axiom widely used in the classical stochastic analysis theory (see Dellacherie and Meyer (1978) [10, Chapter 0.9]). We also rely on the continuum hypothesis, which is needed to apply the aggregation result of Nutz (2012) [26].
2 General model
In order to highlight that our new approach allows to obtain the same results as in [9], we focus here on the exact same model. We believe that this model is sufficiently general to convince the reader of the relevance of our novel approach, and that it could then be extended in a (relatively) straightforward way to solve more complex principal–agent problems.
We fix throughout this paper a finite time horizon $T>0$, and consider the canonical space $\Omega={\cal C}([0,T];\mathbb{R}^{d})$, $d\in\mathbb{N}^{\star}$. The output process $X$ controlled by the agent is originally defined as the canonical process on $\Omega$, i.e.
$X_{t}(x)=x(t),\quad\text{for all }\;x\in\Omega,\;t\in[0,T].$
We denote by $\mathbb{F}=({\cal F}_{t})_{t\geq 0}$ its canonical filtration, defined for all $t\in[0,T]$ by ${\cal F}_{t}=\sigma(X_{s},s\in[0,t])$. Let ${\cal M}$ be the set of all probability measures on $(\Omega,{\cal F}_{T})$, and define ${\cal M}_{\rm loc}$ the subset of ${\cal M}$ such that, for each $\mathbb{P}\in{\cal M}_{\rm loc}$, $X$ is a $(\mathbb{P},\mathbb{F})$–local martingale whose quadratic variation, denoted $\langle X\rangle$, is absolutely continuous in time with respect to the Lebesgue measure. Using the result by Karandikar (1995) [20], the quadratic variation process $\langle X\rangle$ can be defined pathwise, in the sense that there exists an $\mathbb{F}$–progressively measurable process which coincides with the quadratic variation of $X$, $\mathbb{P}$–a.s. for any $\mathbb{P}\in{\cal M}_{\rm loc}$. Moreover, we can introduce the corresponding $\mathbb{F}$–progressively measurable density process $\widehat{\sigma}^{2}$, defined by
$\widehat{\sigma}^{2}_{t}:=\limsup_{\varepsilon\searrow 0}\dfrac{\langle X\rangle_{t}-\langle X\rangle_{t-\varepsilon}}{\varepsilon},\quad t\in[0,T].$ (2.1)
For all $t\in[0,T]$, $\widehat{\sigma}^{2}_{t}\in\mathbb{S}_{+}^{d}$, recalling that $\mathbb{S}_{+}^{d}$ denotes the set of $d\times d$ positive semi-definite symmetric matrices with real entries. We can thus define the corresponding square root process, simply denoted $\widehat{\sigma}$, in the usual way.
2.1 Controlled output process
To represent the agent’s controls on the drift and the volatility of the output process $X$, we let $U$ be a finite-dimensional Euclidean set and consider the following coefficients:
$\lambda:[0,T]\times\Omega\times U\longrightarrow\mathbb{R}^{n},\quad\text{and}\quad\sigma:[0,T]\times\Omega\times U\longrightarrow\mathbb{M}^{d,n}.$
As in [9], we assume that both coefficients are bounded, and that the mappings $\lambda(\cdot,u)$ and $\sigma(\cdot,u)$ are $\mathbb{F}$-optional for any $u\in U$. In particular, recall that such mappings are non-anticipative, in the sense that $\lambda(t,x,u)=\lambda(t,x_{\cdot\wedge t},u)$ and $\sigma(t,x,u)=\sigma(t,x_{\cdot\wedge t},u)$, for all $t\in[0,T]$ and $x\in\Omega:={\cal C}([0,T],\mathbb{R}^{d})$. We will rigorously define later the set ${\cal U}$ of admissible controls, namely appropriate processes taking values in $U$, and consider, for all $\nu\in{\cal U}$, the following stochastic differential equation (SDE for short), driven by an $n$-dimensional Brownian motion $W$, for $n\in\mathbb{N}^{\star}$,
$\mathrm{d}X_{t}=\sigma(t,X,\nu_{t})\big{(}\lambda(t,X,\nu_{t})\mathrm{d}t+\mathrm{d}W_{t}\big{)},\;t\in[0,T].$
To rigorously introduce the weak formulation for the agent’s problem, one may need to enlarge the original canonical space by considering
$\widebar{\Omega}:=\Omega\times{\cal C}([0,T],\mathbb{R}^{n})\times\mathbb{U},$
where $\mathbb{U}$ is the collection of all finite and positive Borel measures on $[0,T]\times U$, whose projection on $[0,T]$ is the Lebesgue measure. In other words, every $q\in\mathbb{\mathbb{U}}$ can be disintegrated as $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$, for an appropriate Borel measurable kernel $q_{t}$, $t\in[0,T]$. The weak formulation requires to consider a subset of $\mathbb{U}$, namely the set $\mathbb{U}_{0}$ of all $q\in\mathbb{U}$ such that the kernel $q_{t}$ is of the form $\delta_{\phi_{t}}(\mathrm{d}u)$ for some Borel function $\phi$, where as usual, $\delta_{\phi_{t}}$ is the Dirac mass at $\phi_{t}$, $t\in[0,T]$. This space is supporting a canonical process $(X,W,\Pi)$ defined by:
$X_{t}(x,\omega,q)$ $:=x(t),\;W_{t}(x,\omega,q):=\omega(t),\;\Pi(x,\omega,q):=q,\;(t,x,\omega,q)\in[0,T]\times\widebar{\Omega}.$
Then, the canonical filtration $\widebar{\mathbb{F}}:=(\widebar{\cal F}_{t})_{t\in[0,T]}$ is defined as the following $\sigma$-algebra,
¯$\widebar{\cal F}_{t}:=\sigma\Big{(}\big{(}X_{s},W_{s},\Delta_{s}(\varphi)\big{)}\text{ s.t. }(s,\varphi)\in[0,t]\times{\cal C}_{b}\big{(}[0,T]\times U,\mathbb{R}\big{)}\Big{)},\;t\in[0,T],$
where ${\cal C}_{b}([0,T]\times U,\mathbb{R})$ is the set of all bounded continuous functions from $[0,T]\times U$ to $\mathbb{R}$, and
$\Delta_{s}(\varphi):=\int_{0}^{s}\int_{U}\varphi(r,u)\Pi(\mathrm{d}r,\mathrm{d}u),\text{ for any }\;(s,\varphi)\in[0,T]\times{\cal C}_{b}([0,T]\times U,\mathbb{R}).$
Finally, let ${\cal C}^{2}_{b}(\mathbb{R}^{k},\mathbb{R})$ for any $k\in\mathbb{N}^{\star}$ denotes the set of bounded twice continuously differentiable functions from $\mathbb{R}^{k}$ to $\mathbb{R}$, whose first and second derivatives are also bounded. For any $(t,\psi)\in[0,T]\times{\cal C}^{2}_{b}(\mathbb{R}^{d+n},\mathbb{R})$, we set
$M_{t}(\psi):=\psi(X_{t},W_{t})-\int_{0}^{t}\int_{U}\bigg{(}$ $\ \widebar{\Lambda}(s,X,u)\cdot\nabla\psi(X_{s},W_{s})+\frac{1}{2}{\rm Tr}\Big{[}\nabla^{2}\psi(X_{s},W_{s})\big{[}\widebar{\Sigma}\widebar{\Sigma}^{\top}\big{]}(s,X,u)\Big{]}\bigg{)}\Pi(\mathrm{d}s,\mathrm{d}u),$ (2.2)
where $\nabla\psi$ and $\nabla^{2}\psi$ denote the gradient vector and the Hessian matrix of $\psi$, respectively, and where $\widebar{\Lambda}$ and $\widebar{\Sigma}$ are respectively the drift vector and the diffusion matrix of the $(d+n)$-dimensional vector process $(X,W)^{\top}$, i.e.
$\widebar{\Lambda}(s,x,u):=\begin{pmatrix}[\sigma\lambda](s,x,u)\\ \mathbf{0}_{n}\end{pmatrix},\;\widebar{\Sigma}(s,x,u):=\begin{pmatrix}\mathbf{0}_{d,d}&\sigma(s,x,u)\\ \mathbf{0}_{n,d}&\mathrm{I}_{n}\\ \end{pmatrix},\;s\in[0,T],\;x\in\Omega,\;u\in U.$
Let $\widebar{\cal M}$ be the set of all probability measures on $(\widebar{\Omega},\widebar{\cal F}_{T})$. We fix some initial conditions, namely $x_{0}\in\mathbb{R}^{d}$ representing the initial value of $X$, and define the following subset of $\widebar{\cal M}$.
Definition 2.1.
The subset $\widebar{\cal P}\subset\widebar{\cal M}$ is composed of all $\widebar{\mathbb{P}}$ such that
  1. $(i)$
    $M(\psi)$ is a (¯$(\widebar{\mathbb{P}},\widebar{\mathbb{F}})$-local martingale on $[0,T]$ for all $\psi\in{\cal C}^{2}_{b}(\mathbb{R}^{d+n},\mathbb{R});$
  2. $(ii)$
    there exists some $w_{0}\in\mathbb{R}^{n}$ such that ¯$\widebar{\mathbb{P}}[(X_{0},W_{0})=(x_{0},w_{0})]=1;$
  3. $(iii)$
    ¯$\widebar{\mathbb{P}}\big{[}\Pi\in\mathbb{U}_{0}]=1$.
[…]
With this in mind, for any $\mathbb{P}\in\widebar{\cal P}$, we denote by $\widebar{\cal U}(\mathbb{P})$ the set of $\widebar{\mathbb{F}}$-predictable and $U$-valued processes $\nu$ such that (2.3) holds, and consider
$\widebar{\cal U}:=\bigcup_{\mathbb{P}\in\widebar{\cal P}}\widebar{\cal U}(\mathbb{P}),$
as well as for any $\nu\in\widebar{\cal U}$, the set of corresponding probability measures $\widebar{\cal P}(\nu):=\{\mathbb{P}\in\widebar{\cal P}:\nu\in\widebar{\cal U}(\mathbb{P})\}$. To avoid technical considerations (see Remark 2.2 below), we further restrict the admissible controls to those that induce a unique weak solution to (2.3), i.e.:
${\cal U}:=\big{\{}\nu\in\widebar{\cal U}:\widebar{\cal P}(\nu)\;\text{is a singleton}\;\{\mathbb{P}^{\nu}\}\big{\}}.$ (2.4)
We thus have a one-to-one correspondence between any admissible control $\nu\in{\cal U}$ and a probability measure $\mathbb{P}\in{\cal P}$, where ${\cal P}$ is defined accordingly:
${\cal P}:=\bigcup_{\nu\in{\cal U}}\widebar{\cal P}(\nu).$
In the following, we will alternatively refer to the unique probability measure in ${\cal P}$ associated with an admissible control $\nu\in{\cal U}$ as $\mathbb{P}^{\nu}$, or the unique admissible control in ${\cal U}$ associated with a probability measure $\mathbb{P}\in{\cal P}$ as $\nu^{\mathbb{P}}$.
Remark 2.2.
The underlying assumption in (2.4) that each admissible effort $\nu\in{\cal U}$ induces a unique weak solution to SDE (2.3) is not enforced in [9] but is considered here to avoid technical considerations regarding the martingale representation property. We refer to Remark 3.3 for further discussions on this point, but wish to justify here that this assumption is not far-fetched. Indeed, it is relatively common in the study of principal–agent problems to at least assume that the martingale representation is valid (see Hernández Santibáñez (2022) [13, Assumption 2.2. and Remark 2.3] for example), and weak uniqueness for SDE (2.3) is a sufficient and more explicit condition for such representation to hold. Furthermore, as one of the main objectives of principal–agent problems is to determine ‘revealing’ contracts, i.e. contracts that reveal the agent’s optimal efforts, it is actually questionable to consider efforts that would induce several weak solutions for SDE (2.3), especially in view of applications.
2.2 The original principal–agent problem
As usual in principal–agent problems, the principal first chooses a contract $\xi$, which represents the compensation paid to the agent at the final time horizon $T$. Naturally, this contract can only be indexed on what is observable by the principal, and should incentivise the agent to appropriately choose a control $\nu\in{\cal U}$, or equivalently a probability $\mathbb{P}\in{\cal P}$. Here, the principal only observes the output process $X$, and therefore any admissible contract $\xi$ should be ${\cal F}_{T}$-measurable, recalling that $\mathbb{F}:=({\cal F}_{t})_{t\in[0,T]}$ represents the canonical filtration of $X$.
The agent’s problem.
Given a compensation $\xi$, the optimisation problem faced by the agent is defined by
$V_{\rm A}(\xi):=\sup_{\mathbb{P}\in{\cal P}}J_{\rm A}(\xi,\mathbb{P}),\;\text{ with }\;J_{\rm A}(\xi,\mathbb{P}):=\mathbb{E}^{\mathbb{P}}\bigg{[}{\cal K}_{\rm A}^{\mathbb{P}}(T)\,\xi-\int_{0}^{T}{\cal K}_{\rm A}^{\mathbb{P}}(s)c\big{(}s,X,\nu_{s}^{\mathbb{P}}\big{)}\mathrm{d}s\bigg{]},\quad\mathbb{P}\in{\cal P},$ (2.5)
where
  1. $(i)$
    the measurable cost function $c:[0,T]\times{\cal C}([0,T],\mathbb{R})\times U\longrightarrow\mathbb{R}$ is such that for any $u\in U$, the map $(t,x)\longmapsto c(t,x,u)$ is $\mathbb{F}$-optional, and satisfies the following integrability condition:
    $\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c\big{(}t,X,\nu^{\mathbb{P}}_{t}\big{)}\big{|}^{p}\mathrm{d}t\bigg{]}<+\infty,\;\text{for some }p>1;$ (2.6)
  2. $(ii)$
    the discount factor ${\cal K}_{\rm A}^{\mathbb{P}}$ is defined for all $\mathbb{P}\in{\cal P}$ by
    ${\cal K}_{\rm A}^{\mathbb{P}}(t):=\exp\bigg{(}-\int_{0}^{t}k_{\rm A}\big{(}s,X,\nu_{s}^{\mathbb{P}}\big{)}\mathrm{d}s\bigg{)},\quad t\in[0,T],$
    where $k_{\rm A}:[0,T]\times{\cal C}([0,T],\mathbb{R})\times U\longrightarrow\mathbb{R}$ is assumed to be measurable, bounded, and such that for any $u\in U$, the map $(t,x)\longmapsto k_{\rm A}(t,x,u)$ is $\mathbb{F}$-optional.
To ensure that the agent’s problem (2.5) is well-defined, we require minimal integrability on the contracts as in [9, Equation (2.5)], namely
$\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\big{[}|\xi|^{p}\big{]}<+\infty,\;\text{ for some }\;p>1.$ (${\rm I}^{p}_{\rm A}$)
In addition, we assume as usual that the agent has a reservation utility level $R_{\rm A}\in\mathbb{R}$, which conditions his acceptance of the contract. We thus denote by $\Xi$ the set of admissible contracts:
[…]
Finally, we denote by ${\cal P}^{\star}(\xi)$ the set of optimal responses of the agent to a contract $\xi\in\Xi$, i.e.
${\cal P}^{\star}(\xi):=\big{\{}\mathbb{P}^{\star}\in{\cal P}\;\text{such that}\;V_{\rm A}(\xi)=J_{\rm A}(\xi,\mathbb{P}^{\star})\big{\}}.$ (2.8)
Remark 2.3.
As mentioned in [9, Remark 2.1], the model and the approach can be extended in a straightforward way to account for risk-aversion in the agent’s utility function. More precisely, in the agent’s expected utility (2.5), $\xi$ can be replaced by $U_{\rm A}(\xi)$ for an appropriate utility function $U_{\rm A}$, modulo the modification of the integrability condition (${\rm I}^{p}_{\rm A}$). As we choose to follow closely the model considered in [9], we do not consider any utility functions here, but we will address the case of CARA—exponential—utility functions through the examples in Section 4.
The principal’s problem.
We continue to follow the framework described in [9] by defining the principal’s optimisation problem as follows:
$V_{\rm P}:=\sup_{\xi\in\Xi}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)}J_{\rm P}(\xi,\mathbb{P}^{\star}),\;\text{ with }\;J_{\rm P}(\xi,\mathbb{P}):=\mathbb{E}^{\mathbb{P}}\Big{[}{\cal K}_{\rm P}(T)\,U_{\rm P}\big{(}\ell(X)-\xi\big{)}\Big{]},\quad(\xi,\mathbb{P})\in\Xi\times{\cal P},$ (2.9)
where
  1. $(i)$
    $U_{\rm P}:\mathbb{R}\longrightarrow\mathbb{R}$ is a non-decreasing and concave utility function;
  2. $(ii)$
    $\ell:\Omega\longrightarrow\mathbb{R}$ is a liquidation function with linear growth;
  3. $(iii)$
    the discount factor ${\cal K}_{\rm P}$ is defined as follows, with $k_{\rm P}:[0,T]\times\Omega\longrightarrow\mathbb{R}$ measurable, bounded and $\mathbb{F}$-optional,
    ${\cal K}_{\rm P}(t):=\exp\bigg{(}-\int_{0}^{t}k_{\rm P}(s,X)\mathrm{d}s\bigg{)},\quad 0\leq t\leq s\leq T.$
Through (2.9), we implicitly assume as in [9] that, if the agent’s optimal response to a given contract $\xi$ is not unique, then the principal is able to choose the most appropriate one. On the contrary, if a contract $\xi\in\Xi$ is such that ${\cal P}^{\star}(\xi)=\varnothing$, then by convention of the supremum, the corresponding expected utility for the principal would be $\infty$- ∞. We can thus assume without loss of generality that any relevant contract $\xi\in\Xi$ should ensure that there exists at least an optimal response $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$, i.e. ${\cal P}^{\star}(\xi)\neq\varnothing$.
For future reference, we summarise this original problem below.
Problem 1 (Original principal–agent problem).
Consider the non-standard stochastic control problem (2.9),
$V_{\rm P}:=\sup_{\xi\in\Xi}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)}J_{\rm P}(\xi,\mathbb{P}^{\star}),\;\text{ with }\;{\cal P}^{\star}(\xi)\neq\varnothing\;\text{ for }\;\xi\in\Xi.$
Usually, 1 cannot be solved directly: the standard approach is to determine an appropriate form of contracts which simplifies both the agent and the principal’s problems, and such that the restriction to this form of contracts is without loss of generality.
Note that the previous setting and assumptions are relatively standard in continuous-time principal–agent problems, and in particular coincide with the framework described in [9]. Indeed, as the goal of this paper is to provide a new approach to solve such principal–agent problems, which in particular requires to show that this approach allows to achieve the same results obtained by Cvitanić, Possamaï, and Touzi (2018) [9], the decision was made to closely follow their original framework. It is nevertheless possible that some assumptions could be relaxed, or at least modified to consider alternative frameworks. For example, in the case of a principal–agent problem with non-continuous output processes (and many agents), one can refer to the recent paper by Hernández Santibáñez (2022) [13] for more general assumptions. It should be noted, however, that this paper focuses on control processes that can be represented through equivalent changes of measures, thus a priori excluding volatility control.
2.3 The existing approach, via 2BSDEs
In this section, we describe briefly and rather informally the general approach developed by Cvitanić, Possamaï, and Touzi (2018) [9], as this will help in comparing it with our new approach, described later in Section 3.
Revealing contracts.
The general approach developed in [9] involves first identifying a class of contracts, offered by the principal, that are revealing, in the sense that the agent’s optimal response to the contract can easily be computed. More precisely, in the framework under consideration here, we first define the appropriate agent’s hamiltonian, for any $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $:=\sup_{u\in U}\;h_{\rm A}(t,x,y,z,\gamma,u),$ (2.10)
[…]
We then introduce the relevant form of contracts, namely $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the process $Y^{y_{0},Z,\Gamma}$ is defined as the solution to an ordinary differential equation (ODE for short) with random coefficients,
$Y_{t}^{y_{0},Z,\Gamma}=y_{0}-\int_{0}^{t}{\cal H}_{\rm A}\big{(}s,X,Y^{y_{0},Z,\Gamma}_{s},Z_{s},\Gamma_{s}\big{)}\mathrm{d}s+\int_{0}^{t}Z_{s}\cdot\mathrm{d}X_{s}+\dfrac{1}{2}\int_{0}^{t}{\rm Tr}\big{[}\Gamma_{s}\mathrm{d}\langle X\rangle_{s}\big{]},\quad\mathbb{P}\textnormal{--a.s.},\;\text{for all }\mathbb{P}\in{\cal P},$ (2.11)
for some processes $(Z,\Gamma)$ satisfying appropriate integrability conditions. More precisely, for any $\mathbb{F}$-predictable $\mathbb{R}^{d}$-valued process $Z$ and any $\mathbb{F}$-optional $\mathbb{R}$-valued process $Y$, we are led to consider the following norms,
$\|Z\|^{p}_{\mathbb{H}^{p}}:=\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\Bigg{[}\bigg{(}\int_{0}^{T}Z_{t}^{\top}\widehat{\sigma}^{2}_{t}Z_{t}\mathrm{d}t\bigg{)}^{p/2}\Bigg{]},\;\text{ and }\;\|Y\|^{p}_{\mathbb{D}^{p}}:=\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\sup_{t\in[0,T]}\big{|}Y_{t}\big{|}^{p}\bigg{]}.$ (2.12)
We can then define the set ${\cal V}$ of appropriate processes $(Z,\Gamma)$ as in [9, Definition 3.2], recalled below.
Definition 2.4.
Let $y_{0}\in\mathbb{R}$. For any $\mathbb{F}$-predictable processes $(Z,\Gamma)$ taking values in $\mathbb{R}^{d}\times\mathbb{M}^{d}$, define the process $Y^{y_{0},Z,\Gamma}$ using (2.11). We will denote $(Z,\Gamma)\in{\cal V}$ if moreover
  1. $(i)$
    $\|Z\|^{p}_{\mathbb{H}^{p}}<\infty$ and $\|Y^{y_{0},Z,\Gamma}\|^{p}_{\mathbb{D}^{p}}<\infty$ for some $p>1$;
  2. $(ii)$
    there exists $\mathbb{P}\in{\cal P}$ such that
    ${\cal H}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t}\big{)}=h_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t},\nu^{\mathbb{P}}_{t}\big{)},\quad\mathrm{d}t\otimes\mathbb{P}\text{--a.e. on }[0,T]\times\Omega.$
Then, the relevant contracts $\xi\in\Xi$ are those that admit the representation $\xi=Y_{T}^{y_{0},Z,\Gamma}$, where the pair $(Z,\Gamma)\in{\cal V}$ as well as the constant $y_{0}\in\mathbb{R}$ in (2.11) will be optimally chosen by the principal, in order to maximise her expected utility, but also to ensure the participation of the agent.
Agent’s optimal response.
Contracts of the form (2.11) are called revealing, as it is straightforward to compute the agent’s best response to such contracts. Indeed, relying on Ito’s formula, it is proved in [9, Proposition 3.3] that, given a contract of the form $\xi=Y_{T}^{y_{0},Z,\Gamma}$, $(y_{0},Z,\Gamma)\in\mathbb{R}\times{\cal V}$, any optimal response $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ is such that
${\cal H}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t}\big{)}=h_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t},\nu^{\mathbb{P}^{\star}}_{t}\big{)},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\textnormal{--a.e. on }[0,T]\times\Omega.$
For later use, we can define, through a classical measurable selection argument, the following function $u^{\star}$,
$u^{\star}(t,x,y,z,\gamma)\in\operatorname*{arg\,max}_{u\in U}h_{\rm A}(t,x,y,z,\gamma,u),\;(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d},$ (2.13)
and remark that the agent’s best response at time $t\in[0,T]$ to a contract $\xi=Y_{T}^{y_{0},Z,\Gamma}$ is therefore given by $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t}^{y_{0},Z,\Gamma},Z_{t},\Gamma_{t}),\;\mathbb{P}^{\star}$–a.s. One can also prove that the value of the agent’s optimal control problem associated to such contract $\xi$ of the form (2.11) boils down to $y_{0}$, namely $V_{\rm A}(Y_{T}^{y_{0},Z,\Gamma})=y_{0}$. Consequently, to ensure the agent’s participation in the contractual agreement, it is clear that the principal must choose $y_{0}\geq R_{\rm A}$.
Principal’s problem.
Under the restriction to contracts of the form (2.11), the principal’s problem boils down to a more standard stochastic control problem of the following form:
$\widetilde{V}_{\rm P}:=\sup_{y_{0}\geq R_{\rm A}}\underline{V}_{\rm P}(y_{0}),\;\text{ with }\;\underline{V}_{\rm P}(y_{0}):=\sup_{(Z,\Gamma)\in{\cal V}}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T}^{y_{0},Z,\Gamma})}J_{\rm P}\big{(}Y_{T}^{y_{0},Z,\Gamma},\mathbb{P}^{\star}\big{)}.$ (2.14)
In particular, for fixed $y_{0}\geq R_{\rm A}$, the value $\underline{V}_{\rm P}(y_{0})$ corresponds to the value of a stochastic control problem with two state variables, $X$ and $Y^{y_{0},Z,\Gamma}$. Notice that, under $\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T}^{y_{0},Z,\Gamma})$, the process $X$ admits the following dynamics,
$\mathrm{d}X_{t}=\sigma\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\Big{(}\lambda\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\mathrm{d}t+\mathrm{d}W_{t}\Big{)},\quad t\in[0,T],$ (2.15)
corresponding to [9, Equation (3.6)]. Using this dynamics for $X$ under the optimal effort, one can also compute the dynamics satisfied by the process $Y^{y_{0},Z,\Gamma}$ defined in (2.11) to obtain for all $t\in[0,T]$
$\mathrm{d}Y_{t}^{y_{0},Z,\Gamma}=$ $\ \bigg{(}Z_{t}\cdot[\sigma\lambda]\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}+\dfrac{1}{2}{\rm Tr}\Big{[}\Gamma_{t}\big{[}\sigma\sigma^{\top}\big{]}\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\Big{]}-{\cal H}_{\rm A}\big{(}t,X,Y^{y_{0},Z,\Gamma}_{t},Z_{t},\Gamma_{t}\big{)}\bigg{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\mathrm{d}W_{t},$
similarly to [9, Equation (3.8)]. Using the definition of the Hamiltonian as well as the fact that the optimal effort is defined as one of its maximiser, we can actually further simplify the previous dynamics to obtain,
$\mathrm{d}Y_{t}^{y_{0},Z,\Gamma}=$ $\ \Big{(}c\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}+Y_{t}^{y_{0},Z,\Gamma}k_{\rm A}\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu_{t}^{\mathbb{P}^{\star}}\big{)}\mathrm{d}W_{t},\quad t\in[0,T].$ (2.16)
Again for future reference, it would be convenient to summarise this restricted problem as follows.
Problem 2 (Restriction to contracts of the form (2.11)).
Consider the standard stochastic control problem (2.14),
$\widetilde{V}_{\rm P}:=\sup_{y_{0}\geq R_{\rm A}}\underline{V}_{\rm P}(y_{0}),\;\text{ with }\;\underline{V}_{\rm P}(y_{0}):=\sup_{(Z,\Gamma)\in{\cal V}}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T})}J_{\rm P}\big{(}Y_{T},\mathbb{P}^{\star}\big{)}.$
where the state variable $(X,Y)$ is solution to the following system of SDEs:
$\mathrm{d}X_{t}$ $=\sigma\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\Big{(}\lambda\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\mathrm{d}t+\mathrm{d}W_{t}\Big{)},\qquad\qquad\qquad\qquad\quad\;\;t\in[0,T],$ (2.17a)
$\mathrm{d}Y_{t}$ $=\Big{(}c\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}+Y_{t}k_{\rm A}\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\star}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (2.17b)
coupled through the agent’s optimal effort $\nu^{\mathbb{P}^{\star}}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$ under $\mathbb{P}^{\star}\in{\cal P}^{\star}(Y_{T})$, where the function $u^{\star}$ is defined as a maximiser of the Hamiltonian through (2.13).
In a Markovian setting, one can solve this stochastic control problem as usual, namely by writing the principal’s Hamiltonian as well as the Hamilton-Jacobi-Bellman (HJB for short) equation satisfied by the value function $\underline{V}_{\rm P}(y_{0})$ (see [9, Theorem 3.9]), and finally optimise it over $y_{0}\geq R_{\rm A}$.
More importantly for our purpose in this paper, one can already remark that the value of 2 should be lower than the value in 1, since considering contracts $\xi$ of the form (2.11) is a restriction of the original set $\Xi$ of admissible contracts. This result can be found in [9, Proposition 3.4], and is a direct consequence of [9, Proposition 3.3] mentioned above. For further reference, this result is stated in the lemma below.
Lemma 2.5.
The original value defined by 1 is greater than the value in 2, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}$.
Optimality of the revealing contracts.
To summarise, when restricting to contracts $\xi\in\Xi$ of the form (2.11), one can first easily determine the agent’s optimal response to such contract, namely by computing the maximiser of his Hamiltonian. Then, still restricting to such revealing contract, the principal’s problem boils down to a more standard stochastic problem, namely 2. These two steps are therefore quite straightforward using standard tools from stochastic control theory. However, the last and crucial step requires to prove that the restriction to such revealing contracts is without loss of generality. More precisely, one need to argue that the value of the original principal’s problem, namely $V_{\rm P}$ recalled in 1, actually coincides with the value when restricting to revealing contracts, i.e. $\widetilde{V}_{\rm P}$ in 2. While it is clear that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, as mentioned in Lemma 2.5 above, the main mathematical challenge is to show the converse inequality, namely $V_{\rm P}\leq\widetilde{V}_{\rm P}$. This result, stated in [9, Theorem 3.6], relies on the sophisticated theory of second-order BSDEs. We deliberately omit to develop further this step here, as it is precisely the part we can avoid with the new approach. It nevertheless seems essential to stress that this is the most difficult stage, which probably prevents a wider application of principal–agent problems with volatility control, but also hampers natural extensions of these problems, as discussed in introduction.
3 A new approach
In this section, we first start by formalising in Section 3.1 a ‘first-best’ alternative problem, in which the principal, in addition to offering a contract, also controls the quadratic variation of the output process, through its density with respect to the Lebesgue measure, denoted by $\Sigma$. We then provide in Section 3.2 a straightforward resolution approach for this alternative problem, relying only on the theory of BSDEs. We will finally prove in Section 3.3 that this reformulation actually leads to the exact same solution compared to the original problem considered in Section 2. In other words, from the principal’s point of view, optimising directly the quadratic variation through its density process $\Sigma$ is strictly equivalent to choosing $\Gamma$ in the contract form (2.11).
3.1 The ‘first-best’ formulation
As already mentioned, the approach developed by Cvitanić, Possamaï, and Touzi (2018) [9] mostly relies on the observation that, if the principal has access to the trajectory of $X$, she can actually compute its quadratic variation $\langle X\rangle$ pathwise. In general, observing the quadratic variation would still prevent the principal from deducing the controls on the volatility, hence the relevance of considering contract of the form (2.11), indexed on $X$ and $\langle X\rangle$ through sensitivity parameters $Z$ and $\Gamma$ to be optimised. The main idea behind our reformulation is to assume further that, since the principal can compute the quadratic variation $\langle X\rangle$, she could actually control it directly. In other words, the idea here is to study an alternative ‘first-best’ formulation of the original principal–agent problem with volatility control, in which the principal enforces a contract $\xi$ as well as a quadratic variation characterised by its density process $\Sigma$. The agent should then optimise his effort in order to maximise his expected utility given such contract $\xi$ as before, but now the admissible set of efforts is restricted to those achieving the quadratic variation imposed by the principal. More precisely, we will consider in the reformulated problem that, given a contract $\xi$ and a quadratic variation density process $\Sigma$, the agent should optimally choose his effort $\nu$ on the drift and the volatility of the output process $X$ so that $\mathrm{d}\langle X\rangle_{t}=\Sigma_{t}\mathrm{d}t$ for all $t\in[0,T]$. Then, the principal’s optimisation problem boils down to choosing the optimal contract $\xi$ and the optimal quadratic variation density process $\Sigma$.
To consider the agent’s problem in this ‘first-best’ reformulation, we shall again fix an admissible contract, namely $\xi\in\Xi$, but also an appropriate process $\Sigma$, representing the density of the quadratic variation with respect to the Lebesgue measure, both would be chosen optimally by the principal, anticipating the agent’s best response. With this in mind, we define the admissible set of processes $\Sigma$ by
${\cal S}:=\big{\{}\Sigma\;\mathbb{F}\text{--progressively measurable process, taking values in}\;\mathbb{S}^{d}_{+}\},$ (3.1)
recalling that $\mathbb{S}_{+}^{d}$ denotes the set of $d\times d$ positive semi-definite symmetric matrices with real entries. For any process $\Sigma\in{\cal S}$, we will denote by $\Sigma^{1/2}$ its $(d,d)$-dimensional symmetric square root, and by $\Sigma^{-1/2}$ its pseudo-inverse.
For fixed $\Sigma\in{\cal S}$, the agent’s admissible efforts are now limited to those that achieve the desired quadratic variation. Mathematically, the set of admissible controls should be modified as follows:
${\cal U}^{\circ}(\Sigma)$ :=$:=\big{\{}\nu\in{\cal U}\;:\;\nu_{t}\in U_{t}^{\circ}(X,\Sigma_{t}),\;t\geq 0\},$
[…]
Accordingly, the set ${\cal P}^{\circ}(\Sigma)$ of admissible probability measures is defined by:
[…]
Remark 3.1.
In this framework, as the density of the quadratic variation is fixed to some process $\Sigma\in{\cal S}$, one can actually write the weak formulation for the agent’s problem by means of Girsanov’s theorem. More precisely, one can originally define the canonical process $X$ through the following SDE,
$X_{t}=x_{0}+\int_{0}^{t}\Sigma_{s}^{1/2}\mathrm{d}W^{\circ}_{s},\quad t\in[0,T],$ (3.3)
where $W^{\circ}$ is a $d$-dimensional Brownian Motion. We can show that starting from a weak solution ¯$\widebar{\mathbb{P}}\in{\cal P}$ to (3.3), one can construct, for any $\nu\in{\cal U}^{\circ}(\Sigma)$, a weak solution $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$ to the original SDE (2.3) through absolute continuous change of measure.
We proceed in two steps. We first relate the previous SDE to the following one:
$X_{t}=x_{0}+\int_{0}^{t}\sigma(s,X,\nu_{s})\mathrm{d}W_{s},\quad t\in[0,T],$ (3.4)
where $W$ is a $n$-dimensional Brownian Motion. For this, we consider an additional $(n-d)$-dimensional Brownian Motion $W^{1}$, potentially defined on an extended probability space, such that $\widebar{W}:=(W^{\circ},W^{1})$ is an $n$-dimensional Brownian Motion. Moreover, denote by ${\cal R}_{n}$ the set of all matrices of rotations of $\mathbb{M}^{n}$, i.e. $R\in\mathbb{M}^{n}$ such that $RR^{\top}=I_{n}$. For any $R\in{\cal R}_{n}$, we will further denote by $R^{\circ}$ the $(d,n)$-dimensional matrix consisting of the $d$ first rows of $R$, and $R^{1}$ the $(n-d,n)$-dimensional matrix consisting of the $n-d$ remaining rows of $R$. For a given control process $\nu\in{\cal U}^{\circ}(\Sigma)$, let $(R_{t})_{t\in[0,T]}$ be an ${\cal R}_{n}$-valued optional process with
$R_{t}^{\circ}:=\Sigma_{t}^{-1/2}\sigma(t,X,\nu_{t}),\quad t\in[0,T],$
where $\Sigma_{t}^{-1/2}$ denotes the pseudo-inverse of $\Sigma_{t}^{1/2}$ for all $t\in[0,T]$, and $(R^{1}_{t})_{t\in[0,T]}$ arbitrary. By Levy’s characterisation of Brownian Motion, the process $W$ defined by $\mathrm{d}W_{t}=R_{t}^{\top}\mathrm{d}\widebar{W}_{t}$ for all $t\in[0,T]$ is also a $\widebar{\mathbb{P}}$–Brownian Motion, and SDE (3.3) can be rewritten as
$X_{t}=x_{0}+\int_{0}^{t}\Sigma_{s}^{1/2}\mathrm{d}W^{\circ}_{s}=x_{0}+\int_{0}^{t}\Sigma_{s}^{1/2}R_{s}^{\circ}\mathrm{d}W_{s}=x_{0}+\int_{0}^{t}\sigma(s,X,\nu_{s})\mathrm{d}W_{s},\quad t\in[0,T].$
In other words, any weak solution $\widebar{\mathbb{P}}$ to SDE (3.3) is also a weak solution to SDE (3.4). Note that the converse also holds, simply by Levy’s characterisation of Brownian Motion.
Consider then the following exponential local martingale
$M^{\nu}_{t}:={\cal E}_{t}\bigg{(}\int_{0}^{\cdot}\lambda(s,X,\nu_{s})\mathrm{d}W_{s}\bigg{)},\quad t\in[0,T],$ (3.5)
where ${\cal E}$ denotes the Doleans–Dade exponential, namely
${\cal E}_{t}(M):=\exp\bigg{(}M_{t}-\dfrac{1}{2}\langle M\rangle_{t}\bigg{)},\quad t\in[0,T],$
for any local martingale $M$. As $\lambda$ is assumed to be bounded, the process $M^{\nu}$ defined above is actually a martingale, and one can thus apply Girsanov’s theorem. In particular, by defining the following change of probability,
$\dfrac{\mathrm{d}\mathbb{P}^{\nu}}{\mathrm{d}\widebar{\mathbb{P}}}\bigg{|}_{{\cal F}_{T}}=M^{\nu}_{T},$
the process $W^{\nu}$ defined by
$W^{\nu}_{\cdot}:=W_{\cdot}-\int_{0}^{\cdot}\lambda(s,X,\nu_{s})\mathrm{d}s$
is a $n$-dimensional $\mathbb{P}^{\nu}$–Brownian Motion, and $X$ admits the following representation:
$X_{t}=x_{0}+\int_{0}^{t}[\sigma\lambda](s,X,\nu_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,X,\nu_{s})\mathrm{d}W^{\nu}_{s},\quad t\in[0,T].$
Therefore, starting from a weak solution $\widebar{\mathbb{P}}$ to (3.4), any admissible control process $\nu\in{\cal U}^{\circ}(\Sigma)$ induces a weak solution $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$ to SDE (2.3). The converse is also true, simply by inverting the above construction.
Informally, the previous remark highlights that, when the quadratic variation is fixed, then the agent’s impact on the output process $X$ can be defined as equivalent change of measures through Girsanov’s theorem. In other words, for fixed $\Sigma\in{\cal S}$, the set ${\cal P}^{\circ}(\Sigma)$ is dominated by $\widebar{\mathbb{P}}$, weak solution to (3.4). Note that $\widebar{\mathbb{P}}$ obviously depends on the original choice of $\Sigma\in{\cal S}$, but for simplicity we decide to omit this dependency in the notations, especially since $\Sigma\in{\cal S}$ will be fixed to solve the agent problem.
Remark 3.2.
The previous remark is also mentioned in [9, Section 4.4]. Indeed, in order to deal with volatility control in the original stochastic control problem for the agent, the idea is to first isolate the control of the quadratic variation, thus considering the following reformulation of the agent’s problem:
$V_{\rm A}(\xi)=\sup_{\Sigma\in{\cal S}}\sup_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma)}J_{\rm A}(\xi,\mathbb{P}).$
For fixed $\Sigma\in{\cal S}$, all probability measures $\mathbb{P}\in{\cal P}^{\circ}(\Sigma)$ are equivalent through Girsanov’s theorem, as mentioned in the previous remark. One can therefore represent the value function of the agent for fixed quadratic variation as a backward stochastic differential equation. Then, since the agent can optimise over admissible quadratic variations, his problem has to be written as the supremum of BSDEs, namely a second-order BSDE. While we will follow the exact same first step, namely to represent the agent’s continuation utility as a BSDE, we will not have to deal with the second step as in our reformulated ‘first-best’ problem, the choice of the quadratic variation is not delegated to the agent, but fixed by the principal.
We can finally formalise the principal–agent problem in this ‘first-best’ scenario. More precisely, given a contract $\xi$ and a quadratic variation density process $\Sigma\in{\cal S}$, the optimisation problem faced by the agent is defined by
$V^{\circ}_{\rm A}(\xi,\Sigma):=\sup_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma)}J_{\rm A}(\xi,\mathbb{P}).$ (3.6)
In parallel to (2.7) and (2.8), for fixed $\Sigma\in{\cal S}$, we define the set $\Xi^{\circ}(\Sigma)$ of admissible contracts as
[…]
and denote by ${\cal P}^{\circ,\star}(\xi,\Sigma)$ the set of agent’s optimal response.
Then, the goal of the principal is to maximise her objective function by choosing the optimal quadratic variation process $\Sigma\in{\cal S}$ as well as the optimal contract $\xi\in\Xi^{\circ}(\Sigma)$, anticipating the agent’s optimal response:
$V_{\rm P}^{\circ}:=\sup_{\Sigma\in{\cal S}}\sup_{\xi\in\Xi^{\circ}(\Sigma)}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\circ,\star}(\xi,\Sigma)}J_{\rm P}(\xi,\mathbb{P}^{\star}).$ (3.8)
Recall that by convention, the supremum over an empty set is equal to $\infty$- ∞. In particular, if the principal suggests a quadratic variation which is not implementable by the agent, then the set ${\cal P}^{\circ,\star}(\xi,\Sigma)$ would be empty, leading to a utility of $\infty$- ∞ for the principal, which cannot be optimal. Therefore, we can assume without loss of generality that any potentially optimal $\Sigma\in{\cal S}$ should be such that ${\cal P}^{\circ}(\Sigma)\neq\varnothing$, or alternatively, ${\cal U}^{\circ}(\Sigma)\neq\varnothing$. In other words, for any $\Sigma\in{\cal S}$, there should exist an admissible effort $\nu\in{\cal U}^{\circ}(\Sigma)$, and thus a corresponding probability measure $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$, such that for all $t\in[0,T]$, $\Sigma_{t}=[\sigma\sigma^{\top}](t,X,\nu_{t}),\;\mathbb{P}^{\nu}$–a.s. This implies in particular that for all $t\in[0,T]$, $\Sigma_{t}$ should take values in $\mathbb{S}_{t}(X)$, defined for all $x\in\Omega$ by
$\mathbb{S}_{t}(x):=\big{\{}S\in\mathbb{S}_{+}^{d}\;:\;S=\big{[}\sigma\sigma^{\top}\big{]}(t,x,u)\;\text{ for some }u\in U\big{\}}.$
Furthermore, recall that the original set ${\cal U}$ of admissible controls is defined such that, for any $\nu\in{\cal U}$, there exists a unique weak solution to (2.3). In particular, the set ${\cal U}^{\circ}(\Sigma)$ of admissible (constrained) controls also satisfies this property. Therefore, under the previous restriction that $\Sigma\in{\cal S}$ should be such that ${\cal U}^{\circ}(\Sigma)\neq\varnothing$, it is clear that weak uniqueness also holds for SDE (3.3). In other words, for fixed $\Sigma\in{\cal S}$, the weak solution $\widebar{\mathbb{P}}$ to SDE (3.3) defined in Remark 3.1 is actually unique, and therefore admits the predictable martingale representation property (see for example Jacod and Shiryaev (2003) [19, Theorem III.4.29]).
Remark 3.3.
Recall that through the original definition of admissible efforts in (2.4), we assume that each $\nu\in{\cal U}$ induces a unique weak solution to (2.3), which implies, as mentioned above, that for fixed $\Sigma\in{\cal S}$, the SDE (3.3) admits a unique weak solution denoted $\widebar{\mathbb{P}}$. This assumption, which is not enforced in [9], is considered here to avoid technical considerations regarding the martingale representation property. Indeed, without weak uniqueness, the solution $\widebar{\mathbb{P}}$ to SDE (3.3) may not admit the predictable martingale representation property. More precisely, the decomposition of any $(\mathbb{F},\widebar{\mathbb{P}})$-local martingale with respect to $X$ would in general involve an extra martingale term, orthogonal to $X$. As this martingale representation property is used to derive the optimal form of contracts, such an additional martingale term would complicate the study. In [9], this additional martingale term can be hidden in the component ‘$K$’ of the solution $(Y,Z,K)$ to the appropriate second-order BSDE, as the set of probability measures considered is naturally saturated (see Possamaï, Tan, and Zhou (2018) [28, Definitions 5.1 and 5.2]). We refer back to Remark 2.2 for a discussion on the reasonableness of this assumption.
Again for future reference, we can summarise this alternative ‘first-best’ problem as follows.
Problem 3.
The ‘first-best’ reformulation of the original principal–agent problem is defined by
$V_{\rm P}^{\circ}:=\sup_{\Sigma\in{\cal S}}\sup_{\xi\in\Xi^{\circ}(\Sigma)}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\circ,\star}(\xi,\Sigma)}J_{\rm P}(\xi,\mathbb{P}^{\star}),$ (3.9)
where ${\cal P}^{\circ,\star}(\xi,\Sigma):=\{\mathbb{P}^{\star}\in{\cal P}^{\circ}(\Sigma)\;:\;V^{\circ}_{\rm A}(\xi,\Sigma)=J_{\rm A}(\xi,\mathbb{P}^{\star})\}\neq\varnothing$ for $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$.
We first provide in Section 3.2 a very standard approach, relying on BSDEs theory, to characterise the optimal form of contracts for this ‘first-best’ alternative problem. The result is given in Theorem 3.9. We then prove the main result of this paper, namely Theorem 3.10, which states the equivalence between 1 and 3, i.e. $V_{\rm P}=V_{\rm P}^{\circ}$. Before proceeding, one can already notice that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem summarised above, namely 3, is greater than the value $V_{\rm P}$ in the original problem (1). This result is stated in the following Lemma 3.4, whose proof mostly relies on the observation that, in 3, the principal has an additional control compared to the original problem.
Lemma 3.4.
The value of 3 is greater than the value of 1, i.e. $V_{\rm P}^{\circ}\geq V_{\rm P}$.
  • Proof of Lemma 3.4.
    Recall that we can assume, without loss of generality, that any potentially optimal contract $\xi\in\Xi$ for 1 should be such that ${\cal P}^{\star}(\xi)\neq\varnothing$, otherwise $V_{\rm P}=-\infty$. One can thus consider $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$, and its corresponding optimal effort process $\nu^{\mathbb{P}^{\star}}\in{\cal U}$, leading to the following representation for the output process $X$:
[…]
    This naturally implies that the quadratic variation of $X$, namely $\langle X\rangle$, observable by the principal, satisfies
[…]
    Note that the process $\Sigma^{\mathbb{P}^{\star}}$ defined above belongs to the set ${\cal S}$ by definition and we have $\nu^{\mathbb{P}^{\star}}\in{\cal U}^{\circ}(\Sigma^{\star})$, or equivalently, $\mathbb{P}^{\star}\in{\cal P}^{\circ}(\Sigma^{\mathbb{P}^{\star}})$. Additionally, by definition of $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$ in (2.8), we have $V_{\rm A}(\xi)=J_{\rm A}(\xi,\mathbb{P}^{\star})\geq R_{\rm A}$, and therefore
    $R_{\rm A}\leq V_{\rm A}(\xi)=J_{\rm A}(\xi,\mathbb{P}^{\star})\leq\sup_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma^{\mathbb{P}^{\star}})}J_{\rm A}(\xi,\mathbb{P})=:V_{\rm A}^{\circ}\big{(}\xi,\Sigma^{\mathbb{P}^{\star}}\big{)},$
    implying in particular that $\xi\in\Xi^{\circ}(\Sigma^{\mathbb{P}^{\star}})$. On the other hand, since we naturally have $V_{\rm A}(\xi)\geq V_{\rm A}^{\circ}(\xi,\Sigma)$ for all $\Sigma\in{\cal S}$ and $\xi\in\Xi$, we can also deduce from the previous inequality that $\mathbb{P}^{\star}\in{\cal P}^{\circ,\star}(\xi,\Sigma^{\mathbb{P}^{\star}})$. To summarise, for any $\xi\in\Xi$ and $\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)$, there exists $\Sigma^{\mathbb{P}^{\star}}\in{\cal S}$ such that $\xi\in\Xi^{\circ}(\Sigma^{\mathbb{P}^{\star}})$ and $\mathbb{P}^{\star}\in{\cal P}^{\circ,\star}(\xi,\Sigma^{\mathbb{P}^{\star}})$, which implies
    $\sup_{\xi\in\Xi}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\star}(\xi)}J_{\rm P}(\xi,\mathbb{P}^{\star})\leq\sup_{\Sigma\in{\cal S}}\sup_{\xi\in\Xi^{\circ}(\Sigma)}\sup_{\mathbb{P}^{\star}\in{\cal P}^{\circ,\star}(\xi,\Sigma)}J_{\rm P}(\xi,\mathbb{P}^{\star})$
    and therefore concludes the proof. ∎
3.2 Solving the ‘first-best’ problem
We demonstrate in this section that the alternative ‘first-best’ problem, namely 3, can be ‘solved’ relying only on the theory of BSDEs. More precisely, one can introduce a relevant form of contracts, which will be proved to be without loss of generality, in order to reformulate the principal’s problem into a more standard stochastic control problem, defined as 4 below. In other words, while 1 is proved to be equivalent to 2 in [9] using the theory of second-order BSDEs, we prove in this section that 3 is equivalent to 4, relying on BSDEs only.
To be consistent with the approach developed in [9], we first introduce the appropriate hamiltonian for the agent when the quadratic variation is fixed, as well as the relevant form of contracts for the reformulated problem (3.8). With this in mind, let $(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}^{d}_{+}$, and consider the following Hamiltonian:
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)$ $:=\sup_{u\in U_{t}^{\circ}(x,S)}\;h^{\circ}_{\rm A}(t,x,y,z,u),$ (3.10)
[…]
recalling that the set $U_{t}^{\circ}(x,S)$ is defined by (3.1).
Remark 3.5.
The Hamiltonian defined above in (3.10) is a simplified version of the original Hamiltonian in (2.10). Indeed, we consider here that the quadratic variation is fixed to some level through its density process $\Sigma$, and in particular, cannot be optimised by the agent. In other words, at each time $t\in[0,T]$, the agent has to choose his effort $\nu_{t}$ taking values in $U$ and satisfying $[\sigma\sigma^{\top}](t,X,\nu_{t})=\Sigma_{t}$. Under this constraint, the ‘$\gamma$ part’ of $h_{\rm A}$ defined in (2.10) becomes a constant, uncontrolled by the agent. Mathematically speaking, we have:
$\sup_{u\in U_{t}^{\circ}(x,S)}h_{\rm A}(t,x,y,z,\gamma,u)$ $=\sup_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}(t,x,y,z,u)+\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}.$
The previous equality also highlights the following relationship between the two Hamiltonians,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)+\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$ (3.11)
In our reformulated problem, because the quadratic variation is originally fixed by the principal, the relevant form of contract will actually be similar to the one considered in principal–agent problems with drift control only. More precisely, we are led to consider here $\xi=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to the following ODE with random coefficients,
$Y_{t}^{y_{0},Z}=y_{0}-\int_{0}^{t}{\cal H}_{\rm A}^{\circ}\big{(}s,X,Y^{y_{0},Z}_{s},Z_{s},\Sigma_{s}\big{)}\mathrm{d}s+\int_{0}^{t}Z_{s}\cdot\mathrm{d}X_{s},\quad\mathbb{P}\textnormal{--a.s.},\;\text{for all }\mathbb{P}\in{\cal P}^{\circ}(\Sigma).$ (3.12)
for some process $Z$ satisfying appropriate integrability conditions. More precisely, in parallel with Definition 2.4, we define the set ${\cal V}^{\circ}(\Sigma)$ of appropriate processes $Z$ as follows.
Definition 3.6.
Let $y_{0}\in\mathbb{R}$ and $\Sigma\in{\cal S}$. For any $\mathbb{F}$–predictable processes $Z$ taking values in $\mathbb{R}^{d}$, define the process $Y^{y_{0},Z}$ using (3.12). We will denote $Z\in{\cal V}^{\circ}(\Sigma)$ if moreover
  1. $(i)$
    $\|Z\|^{p}_{\mathbb{H}^{p}}<\infty$ and $\|Y^{y_{0},Z}\|^{p}_{\mathbb{D}^{p}}<\infty$ for some $p>1;$
  2. $(ii)$
    there exists $\mathbb{P}\in{\cal P}^{\circ}(\Sigma)$ such that
    ${\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}=h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu_{t}^{\mathbb{P}}\big{)},\quad\mathrm{d}t\otimes\mathbb{P}\text{--a.e. on }[0,T]\times\Omega.$ (3.13)
For fixed $\Sigma\in{\cal S}$, any contract defined by $\xi=Y_{T}^{y_{0},Z}$ through (3.12) for some $y_{0}\in\mathbb{R}$ and $Z\in{\cal V}^{\circ}(\Sigma)$ is an ${\cal F}_{T}$-measurable random variable satisfying the integrability condition (${\rm I}^{p}_{\rm A}$). If this contract satisfies in addition the agent’s participation constraint, we conclude that $\xi\in\Xi^{\circ}(\Sigma)$. Conversely, we will prove below in Theorem 3.9 that the restriction to contracts of the form $\xi=Y_{T}^{y_{0},Z}$, for $(y_{0},Z)\in[R_{\rm A},\infty)\times{\cal V}^{\circ}(\Sigma)$, is without loss of generality.
Remark 3.7.
Intuitively, when the density $\Sigma\in{\cal S}$ of the quadratic variation is fixed, the new form of contracts (3.12) can be derived from the general form of contracts (2.11). Indeed, starting from (2.11) under $\mathbb{P}\in{\cal P}^{\circ}(\Sigma)$, and using in particular Remark 3.5 on the Hamiltonian, we derive the following representation,
$Y_{t}^{y_{0},Z,\Gamma}$ $=y_{0}-\int_{0}^{t}\bigg{(}{\cal H}^{\circ}_{\rm A}\big{(}s,X,Y^{y_{0},Z,\Gamma}_{s},Z_{s},\Sigma_{s}\big{)}+\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{s}\Sigma_{s}\big{]}\bigg{)}\mathrm{d}s+\int_{0}^{t}Z_{s}\cdot\mathrm{d}X_{s}+\dfrac{1}{2}\int_{0}^{t}{\rm Tr}\big{[}\Gamma_{s}\Sigma_{s}\big{]}\mathrm{d}s,\;t\in[0,T],$
thus naturally leading to (3.12), since the part of the contract indexed by $\Gamma$ simplifies.
Before stating the main result of this section, we first solve the agent’s problem (3.6), in order to confirm the intuition on the relevant contract form (3.12), already supported by the previous remark. The proof of the following result only relies on the theory of BSDEs, as opposed to 2BSDEs, needed to characterise the agent’s problem originally defined in (2.5) (see [9, Proposition 4.6]).
Proposition 3.8.
Let $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$. There exist a unique ${\cal F}_{0}^{\widebar{\mathbb{P}}+}$-measurable random variable $Y_{0}$, satisfying $\mathbb{E}^{\widebar{\mathbb{P}}}\left[Y_{0}\right]\geq R_{\rm A}$, and a unique $Z\in\mathbb{H}^{p}$ such that $\xi=Y_{T}$, for $Y\in\mathbb{D}^{p}$ defined as the solution to the following ODE:
[…]
Moreover,
  1. $(i)$
    $V_{\rm A}^{\circ}(\xi,\Sigma)=\mathbb{E}^{\widebar{\mathbb{P}}}\left[Y_{0}\right]\geq R_{\rm A};$
  2. $(ii)$
    $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ if and only if ${\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})=h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t}^{\mathbb{P}^{\circ}})$, $\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{--a.e. on }[0,T]\times\Omega$.
  • Proof of Proposition 3.8.
    Step 1. We fix $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$ to consider the following BSDE
[…]
    and argue that there exists a unique solution $(Y,Z)\in\mathbb{D}^{p}\times\mathbb{H}^{p}$ for some $p>1$, to deduce (3.14). For this, we let $p_{1}>1$ and $p_{2}>1$ such that (2.6) and Condition ${\rm I}^{p}_{\rm A}$ hold, respectively. For all $p^{\prime}\in(1,p_{1})$, using Jensen’s inequality and changing the probability from $\widebar{\mathbb{P}}$ to $\mathbb{P}^{\nu}$ using $M^{\nu}$ defined in (3.5), we have
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}M_{T}^{\nu}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{]}$
    $\leq\mathbb{E}^{\mathbb{P}^{\nu}}\Bigg{[}\bigg{(}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p^{\prime}}\mathrm{d}t\bigg{)}^{q_{1}}\Bigg{]}^{\frac{1}{q_{1}}}\mathbb{E}^{\mathbb{P}^{\nu}}\Big{[}\big{|}M_{T}^{\nu}\big{|}^{q_{2}}\Big{]}^{\frac{1}{q_{2}}},$
    where the last inequality comes from Holder’s inequality with $q_{1}:=p_{1}/p^{\prime}>1$ and $q_{2}>1$ its Holder’s conjugate, i.e. such that $\frac{1}{q_{1}}+\frac{1}{q_{2}}=1$. Since $\lambda$ is bounded, there exists a constant $C>0$ such that
    $\mathbb{E}^{\widebar{\mathbb{P}}}\Bigg{[}\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p^{\prime}}\Bigg{]}$ $\leq C\;\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}\inf_{u\in U^{\circ}(\Sigma)}\big{|}c(t,X,u)\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}}\leq C\sup_{\mathbb{P}\in{\cal P}}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}\big{|}c(t,X,\nu_{t})\big{|}^{p_{1}}\mathrm{d}t\bigg{]}^{\frac{1}{q_{1}}},$
    which is finite since (2.6) holds for $p_{1}>1$. By similar arguments, we deduce from the fact that Condition ${\rm I}^{p}_{\rm A}$ is satisfied for $p_{2}>1$ that $\mathbb{E}^{\widebar{\mathbb{P}}}[|\xi|^{p^{\prime\prime}}]<+\infty$, for $p^{\prime\prime}\in(1,p_{2})$. We thus obtain the following integrability condition,
    $\mathbb{E}^{\widebar{\mathbb{P}}}\bigg{[}|\xi|^{p}+\bigg{(}\int_{0}^{T}\big{|}{\cal H}^{\circ}_{\rm A}(t,X,0,0,\Sigma_{t})\big{|}\mathrm{d}t\bigg{)}^{p}\bigg{]}<\infty,\quad p\in(1,p_{1}\wedge p_{2}).$
    Finally, by the boundedness of $\sigma,\lambda$ and $k_{\rm A}$, the Hamiltonian ${\cal H}_{\rm A}^{\circ}$ is uniformly Lipschitz-continuous in $(y,z)$. Since the assumptions of Briand, Delyon, Hu, Pardoux, and Stoica (2003) [6, Theorem 4.2] are satisfied (see also Bouchard, Possamaï, Tan, and Zhou (2018) [5, Theorem 4.1] for a more recent result), and $\widebar{\mathbb{P}}$ satisfies the predictable martingale representation property, we deduce that there exists a unique solution $(Y,Z)\in\mathbb{D}^{p}\times\mathbb{H}^{p}$ to BSDE (3.15), for some $p>1$. In particular, such solution satisfies Definition 3.6 $(i)$ and $Y_{0}$ is a ${\cal F}_{0}^{\widebar{\mathbb{P}}+}$-measurable random variable.
Step 2. We now prove the equality $(i)$, namely $V^{\circ}_{\rm A}(\xi,\Sigma)=\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]$. As a direct consequence, we will obtain that $Z\in\mathbb{H}^{p}$, introduced in Step 1, satisfies (3.13) in Definition 3.6, implying that $(ii)$ is verified. By Step 1, any contract $\xi\in\Xi^{\circ}(\Sigma)$ can be written as
[…]
Writing the dynamics under some arbitrary admissible effort $\nu\in{\cal U}^{\circ}(\Sigma)$, i.e. under $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$, we have:
[…]
Applying Ito’s formula and using the definition of $h^{\circ}_{\rm A}$, one can compute, still under $\mathbb{P}^{\nu}\in{\cal P}^{\circ}(\Sigma)$,
${\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(T)Y_{T}-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)c(t,X,\nu_{t})\mathrm{d}t=Y_{0}$ $-\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t$
$+\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)Z_{t}\cdot\sigma(t,X,\nu_{t})\mathrm{d}W_{t}.$
Using this representation in the agent’s objective function, and noticing that, since $Z\in\mathbb{H}^{p}$ and $k_{\rm A},\sigma$ are bounded, the expectation of the stochastic integral term vanishes, we obtain:
$J_{\rm A}\big{(}\xi,\mathbb{P}^{\nu}\big{)}$ $=\mathbb{E}^{\mathbb{P}^{\nu}}\big{[}Y_{0}\big{]}-\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t\bigg{]}.$
Using the Girsanov’s change of probability highlighted in Remark 3.1, we observe that
$\mathbb{E}^{\mathbb{P}^{\nu}}\big{[}Y_{0}\big{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\big{[}M_{T}^{\nu}Y_{0}\big{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\big{[}M_{0}^{\nu}Y_{0}\big{]}=\mathbb{E}^{\widebar{\mathbb{P}}}\big{[}Y_{0}\big{]},$
since $M^{\nu}$ is both an $(\mathbb{F}^{\widebar{\mathbb{P}}},\widebar{\mathbb{P}})$- and an $(\mathbb{F}^{\widebar{\mathbb{P}}+},\widebar{\mathbb{P}})$-martingale (see Neufeld and Nutz (2014) [25, Proposition 2.2]). We thus have:
$J_{\rm A}\big{(}\xi,\mathbb{P}^{\nu}\big{)}=\mathbb{E}^{\widebar{\mathbb{P}}}\big{[}Y_{0}\big{]}-\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t\bigg{]}.$ (3.16)
On the one hand, it is clear by definition of ${\cal H}^{\circ}_{\rm A}$ and $h^{\circ}_{\rm A}$ that $J_{\rm A}(\xi,\mathbb{P}^{\nu})\leq\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]$, from which follows the first inequality $V^{\circ}_{\rm A}(\xi,\Sigma)\leq\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]$. On the other hand, for all $\nu\in{\cal U}^{\circ}(\Sigma)$, we also have:
$V^{\circ}_{\rm A}(\xi,\Sigma)\geq J_{\rm A}\big{(}\xi,\mathbb{P}^{\nu}\big{)}=\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]-\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}^{\nu}}_{\rm A}(t)\big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t})\big{)}\mathrm{d}t\bigg{]}.$
By definition of the supremum, for all $\varepsilon>0$ and $(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}$, there exists $u^{\varepsilon}\in U_{t}^{\circ}(x,S)$ such that ${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)-h^{\circ}_{\rm A}(t,x,y,z,u^{\varepsilon})\leq\varepsilon$. Replacing in the previous inequality, and using in addition the assumption that $k_{\rm A}$ is bounded, we deduce that there exists a constant $C>0$ such that $V^{\circ}_{\rm A}(\xi,\Sigma)\geq\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]-C\varepsilon$. By arbitrariness of $\varepsilon>0$, we conclude that $V^{\circ}_{\rm A}(\xi,\Sigma)\geq\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]$. To summarise, for arbitrary $\nu\in{\cal U}^{\circ}(\Sigma)$, we have:
$\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]=V^{\circ}_{\rm A}\big{(}\xi,\Sigma\big{)}\geq J_{\rm A}\big{(}\xi,\mathbb{P}^{\nu}\big{)},$
and, since $\xi$ is admissible, the participation constraint must be satisfied, which implies $\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]\geq R_{\rm A}$. This already concludes the proof of $(i)$. Finally, recall that we implicitly assume in 3 that for any $\Sigma\in{\cal S}$ and $\xi\in\Xi^{\circ}(\Sigma)$, there should exist an optimal response for the agent, namely $\mathbb{P}^{\circ}\in{\cal P}^{\circ}(\Sigma)$ such that $V^{\circ}_{\rm A}(\xi,\Sigma)=J_{\rm A}(\xi,\mathbb{P}^{\circ})$. Therefore, looking back at (3.16), we deduce from the existence of $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ that
$\mathbb{E}^{\mathbb{P}^{\circ}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}^{\circ}}_{\rm A}(t)\Big{(}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})-h^{\circ}_{\rm A}\big{(}t,X,Y_{t},Z_{t},\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{)}\mathrm{d}t\bigg{]}=0,$
which implies, since the discount factor is (strictly) positive, that
${\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t},Z_{t},\Sigma_{t}\big{)}=h^{\circ}_{\rm A}\big{(}t,X,Y_{t},Z_{t},\nu_{t}^{\mathbb{P}^{\circ}}\big{)},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{--a.e. on }[0,T]\times\Omega.$
In particular, this completes the proof of $(ii)$, and we can further deduce that $Z$ satisfies (3.13) in Definition 3.6. ∎
Proposition 3.8 provides a BSDE representation for the agent’s dynamic value function, given a quadratic variation density process $\Sigma\in{\cal S}$ and a contract $\xi\in\Xi^{\circ}(\Sigma)$. This already induces a first natural representation for any contract $\xi\in\Xi^{\circ}(\Sigma)$, namely as the terminal value of an appropriate process $Y\in\mathbb{D}^{p}$, defined as the well-posed solution to ODE (3.14). The main difference with the desired form of contracts (3.12) lies in the initial value, which is a random variable $Y_{0}$ in (3.14) but a constant $y_{0}$ in (3.12). The following result ultimately proves that the restriction to contracts of the form $\xi=Y_{T}^{y_{0},Z}$ as in (3.12), for some process $Z\in{\cal V}^{\circ}(\Sigma)$ and a constant $y_{0}\geq R_{\rm A}$, is without loss of generality.
Theorem 3.9.
We have $V_{\rm P}^{\circ}=\widetilde{V}^{\circ}_{\rm P}$, where $\widetilde{V}^{\circ}_{\rm P}$ is defined as follows,
$\widetilde{V}^{\circ}_{\rm P}:=\sup_{y_{0}\geq R_{\rm A}}\underline{V}_{\rm P}^{\circ}(y_{0}),\;\text{ with }\;\underline{V}_{\rm P}^{\circ}(y_{0}):=\sup_{\Sigma\in{\cal S}}\sup_{Z\in{\cal V}^{\circ}(\Sigma)}\sup_{\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y_{T}^{y_{0},Z},\Sigma)}J_{\rm P}\big{(}Y_{T}^{y_{0},Z},\mathbb{P}^{\circ}\big{)},$ (3.17)
with, for $\Sigma\in{\cal S}$, $y_{0}\geq R_{\rm A}$ and $Z\in{\cal V}^{\circ}(\Sigma)$,
  1. $(i)$
    $Y^{y_{0},Z}$ is defined as the solution to ODE (3.12);
  2. $(ii)$
    $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}$ if and only if (3.13) holds for $\mathbb{P}^{\circ}$;
  3. $(iii)$
    $V_{\rm A}^{\circ}\big{(}Y^{y_{0},Z},\Sigma\big{)}=y_{0}\geq R_{\rm A}$.
While it is expected that $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$, the previous result highlights that we actually have equality. In other words, to solve the principal’s problem in the ‘first-best’ reformulation stated in 3, one can restrict the study to contracts of the form (3.12), without loss of generality. Once again, we wish to insist on the fact that the proof of the previous theorem, detailed below, mostly follows the arguments already developed in the proof of Proposition 3.8, and therefore does not rely on the theory of 2BSDEs.
  • Proof of Theorem 3.9.
    Step 1. We first argue that the inequality $V_{\rm P}^{\circ}\geq\widetilde{V}^{\circ}_{\rm P}$ is straightforward, since the principal’s problem corresponding to the value $\widetilde{V}^{\circ}_{\rm P}$ is by definition restricted to contracts of the form (3.12). Indeed, as already mentioned, we clearly have that, for fixed $\Sigma\in{\cal S}$, any contract defined by $\xi=Y_{T}^{y_{0},Z}$ through (3.12) for some $y_{0}\in\mathbb{R}$ and $Z\in{\cal V}^{\circ}(\Sigma)$ is an ${\cal F}_{T}$-measurable random variable, satisfying in addition the integrability condition (${\rm I}^{p}_{\rm A}$) by definition of $Z\in{\cal V}^{\circ}$. To conclude that $\xi\in\Xi^{\circ}(\Sigma)$, it suffices to show that this contract $\xi$ satisfies in addition the agent’s participation constraint, namely $(iii)$ in the theorem. To prove this, we can follow the same arguments as in Step 2 in the proof of Proposition 3.8. More precisely, through similar computations, we obtain
    $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}$ $=\sup_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma)}\mathbb{E}^{\mathbb{P}}\bigg{[}y_{0}-\int_{0}^{T}{\cal K}^{\mathbb{P}}_{\rm A}(t)\Big{(}{\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}-h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu^{\mathbb{P}}_{t}\big{)}\Big{)}\mathrm{d}t\bigg{]}$
    $=y_{0}-\inf_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma)}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}}_{\rm A}(t)\Big{(}{\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}-h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu^{\mathbb{P}}_{t}\big{)}\Big{)}\mathrm{d}t\bigg{]}.$
    First, by definition of ${\cal H}_{\rm A}^{\circ}$ and $h_{\rm A}^{\circ}$, and recalling that the discount factor ${\cal K}^{\mathbb{P}}_{\rm A}$ is (strictly) positive, we clearly have $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}\leq y_{0}$. Then, since $Z\in{\cal V}^{\circ}(\Sigma)$, we have, by Definition 3.6 $(ii)$ that there exists $\mathbb{P}^{\circ}\in{\cal P}^{\circ}(\Sigma)$ such that (3.13) holds, namely
    ${\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}=h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu_{t}^{\mathbb{P}^{\circ}}\big{)},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{--a.e. on }[0,T]\times\Omega.$
    We thus deduce the equality $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}=y_{0}$ for such $\mathbb{P}^{\circ}$, with associated effort $\nu^{\mathbb{P}^{\circ}}$ defined as a maximiser of the Hamiltonian ${\cal H}_{\rm A}^{\circ}$. Finally, since $y_{0}\geq R_{\rm A}$, we conclude that $Y_{T}^{y_{0},Z}\in\Xi^{\circ}(\Sigma)$.
Step 2. We now prove the converse inequality, namely $V_{\rm P}^{\circ}\leq\widetilde{V}^{\circ}_{\rm P}$. Recall that, by Proposition 3.8, any contract $\xi\in\Xi^{\circ}(\Sigma)$ can be written as $\xi=Y_{T}$, where
[…]
with $Y_{0}$ an ${\cal F}_{0}^{\widebar{\mathbb{P}}+}$-measurable random variable and $Z\in\mathbb{H}^{p}$. Moreover, we have $V^{\circ}_{\rm A}(\xi,\Sigma)=\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]\geq R_{\rm A}$, and $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ if and only if ${\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\Sigma_{t})=h^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},\nu_{t}^{\mathbb{P}^{\circ}})$, $\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{--a.e. on }[0,T]\times\Omega$. To obtain the desired contract form (3.12), we need to show that, without loss of generality, the random variable $Y_{0}$ can be replaced by a constant $y_{0}\geq R_{\rm A}$, and verify that we can choose $Z\in{\cal V}^{\circ}(\Sigma)$. With this in mind, let $y_{0}:=\mathbb{E}^{\widebar{\mathbb{P}}}[Y_{0}]\geq R_{\rm A}$ and $Z\in\mathbb{H}^{p}$, to consider the contract defined by $\xi=Y_{T}^{y_{0},Z}$ through (3.12). By Step 1, we already have that
$V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}$ $=y_{0}-\inf_{\mathbb{P}\in{\cal P}^{\circ}(\Sigma)}\mathbb{E}^{\mathbb{P}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}}_{\rm A}(t)\Big{(}{\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}-h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu^{\mathbb{P}}_{t}\big{)}\Big{)}\mathrm{d}t\bigg{]},$ (3.18)
and, in particular, $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}\leq y_{0}$. Using the same arguments as in Step 2 in the proof of Proposition 3.8, we can also prove the converse inequality, and thus conclude that $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}=y_{0}$. Finally, recalling again that ${\cal P}^{\circ,\star}(\xi,\Sigma)\neq\varnothing$, we deduce that $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(\xi,\Sigma)$ if and only if
$\mathbb{E}^{\mathbb{P}^{\circ}}\bigg{[}\int_{0}^{T}{\cal K}^{\mathbb{P}^{\circ}}_{\rm A}(t)\Big{(}{\cal H}^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t}\big{)}-h^{\circ}_{\rm A}\big{(}t,X,Y_{t}^{y_{0},Z},Z_{t},\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{)}\mathrm{d}t\bigg{]}=0,$
which directly implies that (3.13) holds for $\mathbb{P}^{\circ}$, or equivalently that $Z$ satisfies (3.13) in Definition 3.6. To summarise, for a contract $Y_{T}^{y_{0},Z}\in\Xi^{\circ}(\Sigma)$, we have $V^{\circ}_{\rm A}\big{(}Y_{T}^{y_{0},Z},\Sigma\big{)}=y_{0}$, and this value is achieved for the optimal effort $\nu^{\mathbb{P}^{\circ}}$, defined as a maximiser of the Hamiltonian ${\cal H}_{\rm A}^{\circ}$. We conclude that both contracts $Y_{T}$ and $Y_{T}^{y_{0},Z}$ generate the same efforts, and induce the same value for the agent. Therefore, without loss of generality, the principal’s optimisation problem can be restricted to contracts of the form (3.12), namely $\xi:=Y_{T}^{y_{0},Z}$, where the process $Y^{y_{0},Z}$ is defined as the solution to ODE (3.12), for $y_{0}\geq R_{\rm A}$. It remains to verify that $Z\in{\cal V}^{\circ}(\Sigma)$, to conclude that $\widetilde{V}_{\rm P}^{\circ}=V_{\rm P}^{\circ}$. Note that $Z\in\mathbb{H}^{p}$ and satisfies Definition 3.6 $(ii)$, as mentioned above. Finally, following similar arguments as in Step 1 in the proof of Proposition 3.8, one can prove that $Y^{y_{0},Z}\in\mathbb{D}^{p}$, so that Definition 3.6 $(i)$ is satisfied. We conclude that $Z\in{\cal V}^{\circ}(\Sigma)$. ∎
To summarise, Theorem 3.9 allows to conclude that, when the quadratic variation is fixed by the principal through its density $\Sigma\in{\cal S}$, the relevant form of contract is given by $\xi:=Y_{T}^{y_{0},Z}$ where the process $Y^{y_{0},Z}$ is defined by (3.12), for some parameters $y_{0}\geq R_{\rm A}$ and $Z\in{\cal V}^{\circ}(\Sigma)$ optimally chosen by the principal. Moreover, since the agent’s optimal response $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y^{y_{0},Z},\Sigma)$ satisfies (3.13), his optimal effort at time $t\in[0,T]$ can be represented by $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t}^{y_{0},Z},Z_{t},\Sigma_{t})$, where the function $u^{\circ}$ coincides with a maximiser of the Hamiltonian ${\cal H}_{\rm A}^{\circ}$, i.e.
$u^{\circ}(t,x,y,z,S)\in\operatorname*{arg\,max}_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}(t,x,y,z,u),\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}.$ (3.19)
Finally, one can notice that, for fixed $y_{0}\geq R_{\rm A}$, $\underline{V}_{\rm P}^{\circ}(y_{0})$ corresponds to the value of a (more) standard stochastic control problem, with two state variables $X$ and $Y^{y_{0},Z}$, controlled through $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. As done to introduce 2, one can compute the dynamics of these state variables under $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y^{y_{0},Z},\Sigma)$, and conclude that solving 3 is equivalent to solving 4, introduced below.
Problem 4.
Consider the (more) standard stochastic control problem (3.17), namely
$\widetilde{V}^{\circ}_{\rm P}:=\sup_{y_{0}\geq R_{\rm A}}\underline{V}_{\rm P}^{\circ}(y_{0}),\;\text{ with }\;\underline{V}_{\rm P}^{\circ}(y_{0}):=\sup_{\Sigma\in{\cal S}}\sup_{Z\in{\cal V}^{\circ}(\Sigma)}\sup_{\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y_{T},\Sigma)}J_{\rm P}\big{(}\mathbb{P}^{\circ},Y_{T}\big{)}.$
where the couple of state variables $(X,Y)$ is the solution to the following system of SDEs:
$\mathrm{d}X_{t}$ $=\sigma\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{(}\lambda\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\mathrm{d}t+\mathrm{d}W_{t}\Big{)},\qquad\qquad\qquad\qquad\quad\;\;t\in[0,T],$ (3.20a)
$\mathrm{d}Y_{t}$ $=\Big{(}c\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}+Y_{t}k\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\Big{)}\mathrm{d}t+Z_{t}\cdot\sigma\big{(}t,X,\nu^{\mathbb{P}^{\circ}}_{t}\big{)}\mathrm{d}W_{t},\quad t\in[0,T],$ (3.20b)
coupled through the agent’s optimal effort $\nu^{\mathbb{P}^{\circ}}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})$ under $\mathbb{P}^{\circ}\in{\cal P}^{\circ,\star}(Y_{T},\Sigma)$.
One can already notice that 4 stated above is extremely similar to 2. Indeed, in both problems, the principal optimises the same criteria, and, apart from the agent’s optimal effort, the dynamics of the state variables in both problems are identical. The only yet main difference between the two problems is that, in 4, this optimal effort depends (among other variables) on $\Sigma$, while, in 2, it is instead a function of $\Gamma$. We will nevertheless prove, in the next section, that there is a one-to-one correspondence between $\Sigma$ and $\Gamma$ so that optimising one or the other actually lead to the same value for the principal, i.e. $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$.
3.3 Penalisation contracts
The goal of this section is to prove the equivalence between the original problem and its ‘first-best’ reformulation, introduced by 3. As mentioned in the introduction, principal–agent problems with moral hazard are often compared to their ‘first-best’ counterpart, in which the principal can directly impose the agent’s controls. It is usually straightforward to prove that the principal’s value in the first-best scenario is greater than her value in the moral hazard case, since in the first-best case, the principal can directly choose the agent’s control. These two problems are equivalent when the principal actually observes the agent’s controls and can design a contract which strongly penalise him if he deviates from the recommended efforts.
We wish to follow the same type of reasoning here. Notice that we already proved in Lemma 3.4 that the principal’s value $V_{\rm P}^{\circ}$ in the alternative ‘first-best’ problem (3) is greater than the value $V_{\rm P}$ in the original problem (1). Therefore, it remains to design appropriate penalisation contracts to show that, starting from 1, one can actually achieve the value $V_{\rm P}^{\circ}$. This will imply the converse inequality, namely $V_{\rm P}\geq V_{\rm P}^{\circ}$, and thus conclude the proof of the main result of this paper, stated below.
Theorem 3.10.
Solving 3 is equivalent to solving 1, i.e. $V_{\rm P}^{\circ}=V_{\rm P}$.
The idea is to start from the original problem (2.9) and restrict the study to contracts of the form (2.11) introduced in [9], hence considering 2. Then, we only have to prove that this problem is equivalent to 4. Indeed, we already have by Lemma 2.5 that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, and by Theorem 3.9 that $V^{\circ}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$. Therefore, it only remains to prove that $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$ to achieve the desired result, namely $V_{\rm P}\geq\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}=V^{\circ}_{\rm P}$.
Lemma 3.11.
For any fixed $y_{0}\geq R_{\rm A}$, there exists a one-to-one correspondence between $\Sigma\in{\cal S}$, $Z\in{\cal V}^{\circ}(\Sigma)$ in 4 and $(Z,\Gamma)\in{\cal V}$ in 2, so that $\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}$. More precisely, for any $(Z,\Gamma)\in{\cal V}$ and the corresponding state variables $(X,Y)$ with dynamics given by (2.17), the corresponding process $\Sigma\in{\cal S}$ is defined by
$\Sigma_{t}\in\operatorname*{arg\,max}_{S\in\mathbb{S}_{t}(X)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},S)+\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{t}S\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
Conversely, for any $\Sigma\in{\cal S},Z\in{\cal V}^{\circ}(\Sigma)$ and the corresponding state variables $(X,Y)$ with dynamics given by (3.20), the corresponding process $\Gamma$ is defined by
$\Gamma_{t}\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma\Sigma_{t}\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega.$
  • Proof of Lemma 3.11.
    Recall that, in 2, for a given couple $(Z,\Gamma)\in{\cal V}$, the agent’s optimal efforts are given by the maximiser of the Hamiltonian ${\cal H}_{\rm A}$, namely $\nu^{\star}_{t}:=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})$ $t\in[0,T]$, where the function $u^{\star}$ is defined by (2.13). On the other hand, in 4, for $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$, the agent’s optimal efforts are given by the maximiser of the (constrained) Hamiltonian ${\cal H}^{\circ}_{\rm A}$, namely $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})$, $t\in[0,T]$, where the function $u^{\circ}$ is defined by (3.19). Therefore, the first step is to prove that for any $(Z,\Gamma)\in{\cal V}$, there exist $\Sigma\in{\cal S}$ such that $Z\in{\cal V}^{\circ}(\Sigma)$ and satisfying
    $\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})=:\nu^{\star}_{t},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$ (3.21)
    To have a one-to-one correspondence, one need to prove that the converse also holds, namely that for any $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$, there exist $\Gamma$ such that $(Z,\Gamma)\in{\cal V}$ and (3.21) holds $\mathrm{d}t\otimes\mathbb{P}^{\circ}$–a.e. on $[0,T]\times\Omega$.
Step 1. To prove the first step, let $(Z,\Gamma)\in{\cal V}$. Note that, by definition of the function $u^{\star}$, we have:
${\cal H}_{\rm A}(t,x,y,z,\gamma)=h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)},\quad(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}.$
Alternatively, by Remark 3.5, we have for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)+\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$
Recall the definition of the function $u^{\circ}$ in (3.19) as the maximiser of the (constrained) Hamiltonian ${\cal H}^{\circ}_{\rm A}$, namely
$u^{\circ}(t,x,y,z,S)\in\operatorname*{arg\,max}_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}(t,x,y,z,u),\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}.$
Therefore, for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have by definition of the set $U_{t}^{\circ}(x,S)$,
${\cal H}_{\rm A}(t,x,y,z,\gamma)$ $=\sup_{S\in\mathbb{S}_{t}(x)}\bigg{\{}h^{\circ}_{\rm A}(t,x,y,z,u^{\circ}(t,x,z,y,S))+\dfrac{1}{2}{\rm Tr}\big{[}\gamma[\sigma(\sigma)^{\top}](t,x,u^{\circ}(t,x,z,y,S))\big{]}\bigg{\}}$
$=\sup_{S\in\mathbb{S}_{t}(x)}\big{\{}h_{\rm A}(t,x,y,z,\gamma,u^{\circ}(t,x,z,y,S))\big{\}},$
where we also used the definition of $h_{\rm A}$ in (2.10). Hence, for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have
$h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\star}(t,x,y,z,\gamma)\big{)}$ $=\sup_{S\in\mathbb{S}_{t}(x)}\big{\{}h_{\rm A}\big{(}t,x,y,z,\gamma,u^{\circ}(t,x,z,y,S)\big{)}\big{\}},$
and we conclude that there exists a map $S^{\star}$ such that for all $(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d}$, we have $S^{\star}(t,x,y,z,\gamma)\in\mathbb{S}_{t}(x)$ and
$u^{\circ}\big{(}t,x,y,z,S^{\star}(t,x,y,z,\gamma)\big{)}=u^{\star}(t,x,y,z,\gamma).$
In other words, for any $(Z,\Gamma)\in{\cal V}$ and the corresponding state variables $(X,Y)$ with dynamics given by (2.17), one can define a process $\Sigma\in{\cal S}$ by taking
$\Sigma_{t}:=S^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t})\in\operatorname*{arg\,max}_{S\in\mathbb{S}_{t}(X)}\bigg{\{}{\cal H}^{\circ}_{\rm A}(t,X,Y_{t},Z_{t},S)+\dfrac{1}{2}{\rm Tr}\big{[}\Gamma_{t}S\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
It is then straightforward to verify that $Z\in{\cal V}^{\circ}(\Sigma)$ and, by the previous reasoning,
$\nu^{\circ}_{t}:=u^{\circ}(t,X,Y_{t},Z_{t},\Sigma_{t})=u^{\star}(t,X,Y_{t},Z_{t},\Gamma_{t}),\quad\mathrm{d}t\otimes\mathbb{P}^{\star}\text{-a.e. on }[0,T]\times\Omega.$
Step 2. To prove that the converse also holds, let $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. We basically want to show that there exist $\Gamma$ such that $(Z,\Gamma)\in{\cal V}$ and (3.21) holds. For this, we introduce the dual problem of ${\cal H}^{\circ}_{\rm A}$, namely
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S):=\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}},\quad(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}.$ (3.22)
First, using again the link (3.11) between both Hamiltonians ${\cal H}_{\rm A}$ and ${\cal H}^{\circ}_{\rm A}$, it is clear that
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\leq{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]},\quad(t,x,y,z,\gamma)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{M}^{d},\;S\in\mathbb{S}_{t}(x),$
and thus, taking the infimum over $\gamma\in\mathbb{M}^{d}$,
${\cal H}^{\circ}_{\rm A}(t,x,y,z,S)\leq\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S).$ (3.23)
Furthermore, as already noticed in [9, Remark 3.5], the Hamiltonian ${\cal H}_{\rm A}$ is convex in $\gamma\in\mathbb{M}^{d}$, implying here that the dual problem introduced above by (3.22) is a convex minimisation problem. Moreover, since there exists an optimal solution for the primal problem and the dual variable is unconstrained, there exists an optimal solution for the dual problem, i.e. a minimiser $\gamma^{\star}:=g^{\star}(t,x,y,z,S)\in\mathbb{M}^{d}$. By definition of $u^{\star}$ and $h_{\rm A}$, we thus have
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=\inf_{\gamma\in\mathbb{M}^{d}}\bigg{\{}h_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma)\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}=h_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}-\dfrac{1}{2}{\rm Tr}\big{[}\gamma^{\star}S\big{]},$
for all $(t,x,y,z,S)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{S}_{+}^{d}$. Moreover, since the dual problem is convex in $\gamma$, we have
$[\sigma\sigma^{\top}](t,x,u^{\star}(t,x,y,z,\gamma))-S\in 2\partial_{\gamma}\bigg{(}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{)},$
where here, $\partial_{\gamma}$ denotes the subgradient with respect to $\gamma$. Since $\gamma^{\star}$ is a minimiser, we necessarily have
$[\sigma(\sigma)^{\top}]\big{(}t,x,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}=S,$ (3.24)
for all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$. We thus deduce, using the definitions of $h^{\circ}_{\rm A}$ and ${\cal H}_{\rm A}^{\circ}$ that
~$\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)}\leq\sup_{u\in U_{t}^{\circ}(x,S)}h^{\circ}_{\rm A}\big{(}t,x,y,z,u\big{)}={\cal H}^{\circ}_{\rm A}(t,x,y,z,S),$
since $u^{\star}(t,x,y,z,\gamma^{\star})\in U^{\circ}_{t}(x,S)$ by (3.24). Combining with (3.23), we conclude that
$h^{\circ}_{\rm A}\big{(}t,x,y,z,S,u^{\circ}(t,x,y,z,S)\big{)}={\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=\widetilde{\cal H}^{\circ}_{\rm A}(t,x,y,z,S)=h^{\circ}_{\rm A}\big{(}t,x,y,z,u^{\star}(t,x,y,z,\gamma^{\star})\big{)},$
for all $(t,x,y,z)\in[0,T]\times\Omega\times\mathbb{R}\times\mathbb{R}^{d}$ and $S\in\mathbb{S}_{t}(x)$, implying in particular
$u^{\circ}(t,x,y,z,S)=u^{\star}(t,x,y,z,\gamma^{\star}),\;\text{ for }\;\gamma^{\star}\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,x,y,z,\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma S\big{]}\bigg{\}}.$
In other words, for any $\Sigma\in{\cal S},Z\in{\cal V}^{\circ}(\Sigma)$ and the corresponding state variables $(X,Y)$ with dynamics given by (3.20), one can define $\Gamma$ as follows,
$\Gamma_{t}:=g^{\star}(t,X,Y_{t},Z_{t},\Sigma_{t})\in\operatorname*{arg\,min}_{\gamma\in\mathbb{M}^{d}}\bigg{\{}{\cal H}_{\rm A}(t,X,Y_{t},Z_{t},\gamma)-\dfrac{1}{2}{\rm Tr}\big{[}\gamma\Sigma_{t}\big{]}\bigg{\}},\quad\mathrm{d}t\otimes\mathbb{P}^{\circ}\text{-a.e. on }[0,T]\times\Omega.$
Moreover, the integrability conditions in Definition 2.4 follows from the integrability conditions in Definition 3.6, implying that $(Z,\Gamma)\in{\cal V}$, and by the previous reasoning, the desired equality of the optimal efforts. ∎
To summarise, using Lemma 2.5, Lemma 3.11, Theorem 3.9 and finally Lemma 3.4, we obtain the following relationship for the values of the various problems,
$V_{\rm P}\geq\widetilde{V}_{\rm P}=\widetilde{V}^{\circ}_{\rm P}=V^{\circ}_{\rm P}\geq V_{\rm P},$
which directly implies equality of all the values, thus concluding the proof of Theorem 3.10. In other words, the four problems outlined are in fact equivalent. While 1 and 3 cannot generally be solved directly, 2 and 4 correspond to more standard stochastic control problems, and should therefore be easier to solve, through standard techniques relying on Hamilton-Jacobi-Bellman equations. We illustrate in the following section the link between 2 and 4 and their solutions through a couple of toy examples with explicit solutions. In particular, these examples will allow to better highlight the link between the controls $\Gamma$ in 2 and $\Sigma$ in 4, which is not quite explicit in the above result Lemma 3.11, because of the general formulation.
4 Illustrative Examples
In this section, we present two illustrative examples, in order to highlight the link between 2 and 4, in particular the one-to-one correspondence between the contract’s sensitivity parameter $\Gamma$ in 2 and the quadratic variation density process $\Sigma$ in 4. The introduction and resolution of these two examples is rather informal, we refer to the previous sections for a more rigorous formulation and the theoretical results.
4.1 A first-best case
We first consider a very simple example, in which the agent only controls the volatility of the output process. More precisely, we let
$X_{t}:=x_{0}+\int_{0}^{t}\nu_{s}\mathrm{d}W_{s},\;t\in[0,T],\;x_{0}\in\mathbb{R},$
where $W$ here is a one-dimensional Brownian Motion, and $\nu$ is an ‘admissible’ control chosen by the agent, namely satisfying appropriate integrability conditions and taking values in $[0,1]$. We further assume that, given a contract $\xi$, the agent wishes to maximise the following expected utility,
$J_{\rm A}(\xi,\nu):=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}U_{\rm A}\bigg{(}\xi-\dfrac{1}{2}\int_{0}^{T}\dfrac{1}{|\nu_{t}|^{2}}\mathrm{d}t\bigg{)}\bigg{]},\;\text{ with }\;U_{\rm A}(x)=-\mathrm{e}^{-\gamma_{\rm A}x},\;x\in\mathbb{R},\;\gamma_{\rm A}>0.$
Similarly, the principal’s expected utility to be maximised is defined by
$J_{\rm P}(\xi,\nu):=\mathbb{E}^{\mathbb{P}^{\nu}}\big{[}U_{\rm P}\big{(}X_{T}-\xi-h\langle X\rangle_{T}\big{)}\big{]},\;\text{ with }\;U_{\rm P}(x)=-\mathrm{e}^{-\gamma_{\rm P}x},\;x\in\mathbb{R},\;\gamma_{\rm P}>0,\;h>0.$
We first recall that, since the principal observes $X$ in continuous time, she can also compute the associated quadratic variation process, namely $\langle X\rangle$. Here, such process follows the dynamics $\mathrm{d}\langle X\rangle_{t}=\nu_{t}^{2}\mathrm{d}t$, $t\in[0,T]$. In other words, the quadratic variation density process, denoted $\Sigma$ throughout the paper, satisfies $\Sigma_{t}=\nu_{t}^{2}$, $t\in[0,T]$, and since $\nu$ takes values in $[0,1]$, it is already clear that, by observing $\Sigma$ in continuous-time, the principal can deduce the agent’s effort $\nu_{t}=\sqrt{\Sigma_{t}}$, $t\in[0,T]$. This example therefore illustrates a first-best scenario, since from her observation, the principal can actually deduce the agent’s effort.
If one follows the approach developed by [9], the optimal form of contract is the following
$\xi=Y_{T}^{y_{0},Z,\Gamma}=U_{\rm A}^{-1}(y_{0})-\int_{0}^{T}{\cal H}_{\rm A}(\Gamma_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\mathrm{d}X_{t}+\dfrac{1}{2}\int_{0}^{T}\big{(}\Gamma_{t}+\gamma_{\rm A}|Z_{t}|^{2}\big{)}\mathrm{d}\langle X\rangle_{t},$
for $y_{0}\in\mathbb{R}$, $(Z,\Gamma)\in{\cal V}$, both taking values in $\mathbb{R}$, and where the agent’s Hamiltonian is defined by
${\cal H}_{\rm A}(\gamma):=\dfrac{1}{2}\sup_{u\in[0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}},\;\gamma\in\mathbb{R}.$
The maximiser of the previous Hamiltonian is given by
$u^{\star}(\gamma):=\begin{cases}(-\gamma)^{-1/4}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1,\end{cases}$
meaning that the agent’s optimal response to the contract $\xi$ is given by $\nu^{\star}_{t}=u^{\star}(\Gamma_{t})$, $t\in[0,T]$. Moreover, given such contract, one can easily compute that the agent’s value boils down to $y_{0}$, and therefore the principal should choose $y_{0}\geq R_{\rm A}$, if $R_{\rm A}$ represents the agent’s reservation utility. Computing the dynamics of $X$ and $Y^{y_{0},Z,\Gamma}$ under the optimal effort, one obtains
$\mathrm{d}X_{t}$ $=u^{\star}(\Gamma_{t})\mathrm{d}W_{t},\text{ and }\;\mathrm{d}Y_{t}^{y_{0},Z,\Gamma}=\dfrac{1}{2}\big{(}|u^{\star}(\Gamma_{t})|^{-2}+\gamma_{\rm A}|Z_{t}|^{2}|u^{\star}(\Gamma_{t})|^{2}\big{)}\mathrm{d}t+Z_{t}u^{\star}(\Gamma_{t})\mathrm{d}W_{t},\quad t\in[0,T].$
To reformulate the principal’s problem in a more simple way, one can actually define a new state variable, namely $L_{t}:=X_{t}-Y_{t}^{y_{0},Z,\Gamma}-h\langle X\rangle_{t}$, $t\in[0,T]$, so that the principal’s problem becomes
$\ \sup_{y_{0}\geq R_{\rm A}}\widetilde{V}_{\rm P}(y_{0}),\;\text{where }\;\widetilde{V}_{\rm P}(y_{0})=\sup_{(Z,\Gamma)\in{\cal V}}\mathbb{E}^{\mathbb{P}^{\nu^{\star}(\Gamma)}}\big{[}U_{\rm P}(L_{T})\big{]},$ (4.1)
with $\ \mathrm{d}L_{t}=-\dfrac{1}{2}\big{(}|u^{\star}(\Gamma_{t})|^{-2}+\gamma_{\rm A}|Z_{t}|^{2}|u^{\star}(\Gamma_{t})|^{2}+2h|u^{\star}(\Gamma_{t})|^{2}\big{)}\mathrm{d}t+u^{\star}(\Gamma_{t})(1-Z_{t})\mathrm{d}W_{t},\;t\in[0,T],$
and with initial condition $L_{0}=x_{0}-U_{\rm A}^{-1}(y_{0})$. Note that for fixed $y_{0}\geq R_{\rm A}$, the previous optimisation problem is a standard stochastic control problem with a one-dimensional state variable $L$ controlled through the parameters $(Z,\Gamma)\in{\cal V}$. For the sake of completeness, the solution is provided at the end of this section, see Remark 4.1.
If we follow our approach, we first need to fix the quadratic variation density process $\Sigma\in{\cal S}$. Note here that, since $\Sigma_{t}=|\nu_{t}|^{2}$, $t\in[0,T]$, and $\nu$ takes values in $[0,1]$, $\Sigma$ is also restricted to take values in $[0,1]$. The (constrained) Hamiltonian ${\cal H}_{\rm A}^{\circ}$ is defined by
${\cal H}_{\rm A}^{\circ}(S)=\dfrac{1}{2}\sup_{u\in[0,1]:u^{2}=S}\big{\{}-u^{-2}\big{\}}=-\frac{1}{2S},\;\text{ for }\;u^{\circ}(S):=\sqrt{S},\;S\in[0,1],$
and for $\Sigma\in{\cal S}$, the relevant form of contract is given by
$\xi^{\circ}=Y_{T}^{y_{0},Z}=U_{\rm A}^{-1}(y_{0})-\int_{0}^{T}{\cal H}^{\circ}_{\rm A}(\Sigma_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\mathrm{d}X_{t}+\dfrac{1}{2}\gamma_{\rm A}\int_{0}^{T}|Z_{t}|^{2}\mathrm{d}\langle X\rangle_{t},$
for $y_{0}\in\mathbb{R}$ and $Z\in{\cal V}^{\circ}(\Sigma)$. For any $\Sigma\in{\cal S}$ and $\xi^{\circ}$ of the previous form, the agent’s best response is given by the maximiser of the (constrained) Hamiltonian ${\cal H}_{\rm A}^{\circ}$, namely $\nu^{\circ}_{t}=u^{\circ}(\Sigma_{t})=\sqrt{\Sigma_{t}}$, $t\in[0,T]$, which allows to achieve the best value $y_{0}\geq R_{\rm A}$. As before, we can compute the dynamics of $X$ and $Y^{y_{0},Z}$ under the optimal effort to obtain
$\mathrm{d}X_{t}$ $=u^{\circ}(\Sigma_{t})\mathrm{d}W_{t},\text{ and }\;\mathrm{d}Y_{t}^{y_{0},Z}=\dfrac{1}{2}\big{(}|u^{\circ}(\Sigma_{t})|^{-2}+\gamma_{\rm A}|Z_{t}|^{2}|u^{\circ}(\Sigma_{t})|^{2}\big{)}\mathrm{d}t+Z_{t}u^{\circ}(\Sigma_{t})\mathrm{d}W_{t},\quad t\in[0,T].$
Defining $L^{\circ}_{t}:=X_{t}-Y_{t}^{y_{0},Z}-h\langle X\rangle_{t}$, $t\in[0,T]$, the principal’s problem becomes a standard stochastic control problem with a one-dimensional state variable $L^{\circ}$, controlled through $\Sigma\in{\cal S}$ and $Z\in{\cal V}^{\circ}(\Sigma)$:
$\ \sup_{y_{0}\geq R_{\rm A}}\widetilde{V}_{\rm P}(y_{0}),\;\text{where }\;\widetilde{V}_{\rm P}(y_{0})=\sup_{\Sigma\in{\cal S}}\sup_{Z\in{\cal V}^{\circ}(\Sigma)}\mathbb{E}^{\mathbb{P}^{\nu^{\circ}(\Sigma)}}\big{[}U_{\rm P}(L^{\circ}_{T})\big{]},$ (4.2)
with $\ \mathrm{d}L^{\circ}_{t}=-\dfrac{1}{2}\big{(}|u^{\circ}(\Sigma_{t})|^{-2}+\gamma_{\rm A}|Z_{t}|^{2}|u^{\circ}(\Sigma_{t})|^{2}+2h\big{|}u^{\circ}(\Sigma_{t})\big{|}^{2}\big{)}\mathrm{d}t+u^{\circ}(\Sigma_{t})(1-Z_{t})\mathrm{d}W_{t},\;t\in[0,T],$
with initial condition $L_{0}=x_{0}-U_{\rm A}^{-1}(y_{0})$.
Comparing (4.1) and (4.2), we clearly need to have equality between the optimal controls to ensure that both problems lead to the same results. This necessitates $u^{\circ}(S)=u^{\star}(\gamma)$, for $(\gamma,S)\in\mathbb{R}\times[0,1]$, and we therefore intuit the following one-to-one correspondence:
$S=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
This is in line with the correspondence given by Lemma 3.11 since, for $\gamma\in\mathbb{R}$, we clearly have
$\operatorname*{arg\,max}_{S\in[0,1]}\bigg{\{}{\cal H}^{\circ}_{\rm A}(S)+\dfrac{1}{2}\gamma S\bigg{\}}=\operatorname*{arg\,max}_{S\in[0,1]}\big{\{}-S^{-1}+\gamma S\big{\}}=\begin{cases}|\gamma|^{-1/2}&\text{if }\gamma<-1,\\ 1&\text{if }\gamma\geq-1.\end{cases}$
On the other hand, for $S\in[0,1]$ and $\gamma\in\mathbb{R}$,
${\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S=\dfrac{1}{2}\sup_{u\in[0,1]}\big{\{}\gamma u^{2}-u^{-2}\big{\}}-\dfrac{1}{2}\gamma S=\begin{cases}-|\gamma|^{1/2}-\dfrac{1}{2}\gamma S&\text{if }\gamma<-1,\\[5.0pt] \dfrac{1}{2}\big{(}\gamma-1-\gamma S\big{)}&\text{if }\gamma\geq-1.\end{cases}$
implying since $S\in[0,1]$ that
$\operatorname*{arg\,min}_{\gamma\in\mathbb{R}}\bigg{\{}{\cal H}_{\rm A}(\gamma)-\dfrac{1}{2}\gamma S\bigg{\}}=-S^{-2}.$
Remark 4.1.
To solve (4.1) or (4.2), one can compute the principal’s Hamiltonian and write the corresponding HJB equation. Recall that both problems are equivalent, but here, (4.2) may be more straightforward to solve compared to (4.1), where one would need to distinguish cases depending on the value of $\Gamma$. Through standard techniques, one obtains the following values and efforts for the agent and the principal, respectively,
$V_{\rm A}(\xi)$ $=R_{\rm A},\;\text{ for }\;\nu^{\star}\equiv 1\wedge\big{(}2h+\gamma_{\rm A}\overline{\gamma}\big{)}^{-1/4},\;\text{ with }\;\overline{\gamma}:=\dfrac{\gamma_{\rm P}}{\gamma_{\rm P}+\gamma_{\rm A}};$
[…]
One can verify that solving (4.1) lead to the exact same results, but with the optimal sensitivity parameter
$\Gamma^{\star}\equiv-\big{|}\Sigma^{\star}\big{|}^{-2}\equiv-\big{(}1\vee\big{(}2h+\gamma_{\rm A}\overline{\gamma}\big{)}\big{)}.$
4.2 A more intricate example
We now consider a model inspired by the demand-response problem introduced and solved by Aïd, Possamaï, and Touzi (2022) [1]. In this model [^2] We will only briefly describe the model here and refer to [1] for further discussion and motivation regarding the problem under consideration., the agent can impact, through an admissible control $\nu:=(\alpha,\beta)\in{\cal U}$, both the mean and the volatility of his electricity consumption for various specific usages, such as heating, air conditioning, etc. His deviation with respect to his original consumption $x_{0}\in\mathbb{R}_{+}$ is denoted $X$, and satisfies:
$X_{t}=x_{0}-\int_{0}^{t}\alpha_{s}\cdot\mathbf{1}_{d}\mathrm{d}s+\int_{0}^{t}\sigma(\beta_{s})\mathrm{d}W_{s},\quad t\in[0,T],$
where $W$ is a $d$-dimensional Brownian Motion. The volatility function $\sigma$ in the previous dynamics is defined by
$\sigma(b):=(\sigma_{1}\sqrt{b_{1}},\dots,\sigma_{d}\sqrt{b_{d}}),\;b:=(b_{1},\dots,b_{d})\in[0,1]^{d},\;\text{ for }\;\sigma_{k}>0,\;k\in\{1,\dots,d\}.$
As usual, it is assumed that the principal observes the output process $X$ in continuous time, and therefore its quadratic variation. However, contrary to the previous example, these observations are not sufficient to deduce the agent’s effective effort.
Given an admissible contract $\xi$, the agent’s expected utility to be maximised is defined as follows,
$J_{\rm A}(\xi,\nu):=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}U_{\rm A}\bigg{(}\xi+\int_{0}^{T}\big{(}\kappa X_{t}-c(\nu_{t})\big{)}\mathrm{d}t\bigg{)}\bigg{]},\;\text{ with }\;U_{\rm A}(x)=-\mathrm{e}^{-\gamma_{\rm A}x},\;x\in\mathbb{R},\;\gamma_{\rm A}>0,\;\kappa>0,$
where the cost of effort is given, for any $u:=(a,b)\in A\times B$ with $A:=\mathbb{R}_{+}^{d}$ and $B:=\mathbb{R}_{+}^{d}$, by
$c(u):=c_{1}(a)+\dfrac{1}{2}c_{2}(b),\;\text{ with }\;c_{1}(a):=\dfrac{1}{2}\sum_{k=1}^{d}\frac{a_{k}^{2}}{\mu_{k}}\;\text{ and }\;c_{2}(b):=\sum_{k=1}^{d}\dfrac{\sigma_{k}^{2}}{\lambda_{k}b_{k}},$
for some parameters $\lambda_{k}>0$ and $\mu_{k}>0$, $k\in\{1,\dots,d\}$. [^3] Note that here, for computational convenience and illustration purposes, we slightly simplify the model, in particular the sets $A$ and $B$. Indeed, in the original model, these sets are bounded, which will drastically complicates the agent’s problem in our approach. The principal, who represents an electricity producer/supplier, wishes to incentivise the consumer to decrease the mean and the volatility of his consumption during the contracting period, in order to maximise his expected utility, defined as follows:
$J_{\rm P}(\xi,\nu):=\mathbb{E}^{\mathbb{P}^{\nu}}\bigg{[}U_{\rm P}\bigg{(}-\xi-\theta\int_{0}^{T}X_{t}\mathrm{d}t-\dfrac{1}{2}h\langle X\rangle_{T}\bigg{)}\bigg{]},\;\text{ with }\;U_{\rm P}(x)=-\mathrm{e}^{-\gamma_{\rm P}x},\;x\in\mathbb{R},\;\gamma_{\rm P}>0,\;h>0,\;\theta>0.$
If we follow the approach developed in [9], and used in [1], the optimal form of contract is identical to the one introduced for the previous example, namely
$\xi=Y_{T}^{y_{0},Z,\Gamma}=U_{\rm A}^{-1}(y_{0})-\int_{0}^{T}{\cal H}_{\rm A}(X_{t},Z_{t},\Gamma_{t})\mathrm{d}t+\int_{0}^{T}Z_{t}\mathrm{d}X_{t}+\dfrac{1}{2}\int_{0}^{T}\big{(}\Gamma_{t}+\gamma_{\rm A}|Z_{t}|^{2}\big{)}\mathrm{d}\langle X\rangle_{t},$
for $y_{0}\in\mathbb{R}$, $(Z,\Gamma)\in{\cal V}$, both taking values in $\mathbb{R}$, but the agent’s Hamiltonian is now defined for $(x,z,\gamma)\in\mathbb{R}^{3}$ by
${\cal H}_{\rm A}(x,z,\gamma)$ $=\sup_{(a,b)\in A\times B}\bigg{\{}-a\cdot\mathbf{1}_{d}z+\dfrac{1}{2}\gamma\sigma(b)\sigma^{\top}(b)+\kappa x-c_{1}(a)-\dfrac{1}{2}c_{2}(b)\bigg{\}}$
$=\kappa x-\inf_{a\in A}\big{\{}a\cdot\mathbf{1}_{d}z+c_{1}(a)\big{\}}-\dfrac{1}{2}\inf_{b\in B}\big{\{}c_{2}(b)-\gamma\sigma(b)\sigma^{\top}(b)\big{\}}.$
Computing the optimiser of the previous Hamiltonian will determine the optimal response to the contract $\xi$ indexed by $(Z,\Gamma)\in{\cal V}$, i.e. for all $k=1,\dots,d$,
$a^{\star}_{k}(z)=-\mu_{k}z\vee 0,\;\text{ and }\;b^{\star}_{k}(\gamma)=(-\lambda_{k}\gamma)^{-1/2},\;\gamma<0.$
We can compute the dynamics of $X$ and $Y^{y_{0},Z,\Gamma}$ under the optimal effort, and rewrite the more standard stochastic control problem of the principal as follows
$\ \sup_{y_{0}\geq R_{\rm A}}\widetilde{V}_{\rm P}(y_{0}),\;\text{where }\;\widetilde{V}_{\rm P}(y_{0})=\sup_{(Z,\Gamma)\in{\cal V}}\mathbb{E}^{\mathbb{P}^{\nu^{\star}(Z,\Gamma)}}\bigg{[}U_{\rm P}\bigg{(}-Y^{y_{0},Z,\Gamma}_{T}-\theta\int_{0}^{T}X_{t}\mathrm{d}t-\dfrac{1}{2}h\langle X\rangle_{T}\bigg{)}\bigg{]},$ (4.3)
with state variables’ dynamics
$\mathrm{d}X_{t}$ $=-\widebar{\mu}(Z_{t}\wedge 0)\mathrm{d}t+(-\Gamma_{t})^{-1/4}\sum_{k=1}^{d}\dfrac{\sigma_{k}}{\lambda_{k}^{1/4}}\mathrm{d}W_{t}^{k},$
$\mathrm{d}Y^{y_{0},Z,\Gamma}_{t}$ $=\dfrac{1}{2}\bigg{(}-2\kappa X_{t}+\widebar{\mu}(Z_{t}\wedge 0)^{2}+\widebar{\sigma}\bigg{(}\sqrt{-\Gamma_{t}}+\gamma_{\rm A}\dfrac{|Z_{t}|^{2}}{\sqrt{-\Gamma_{t}}}\bigg{)}\bigg{)}\mathrm{d}t+\dfrac{Z_{t}}{(-\Gamma_{t})^{1/4}}\sum_{k=1}^{d}\dfrac{\sigma_{k}}{\lambda_{k}^{1/4}}\mathrm{d}W_{t}^{k},$
for $t\in[0,T]$, with initial conditions $X_{0}=x_{0}$ and $Y^{y_{0},Z,\Gamma}_{0}=U_{\rm A}^{-1}(y_{0})$, and where we denoted
$\widebar{\mu}:=\sum_{k=1}^{d}\mu_{k},\;\text{ and }\;\widebar{\sigma}:=\sum_{k=1}^{d}\dfrac{\sigma_{k}^{2}}{\sqrt{\lambda_{k}}}.$
Proceeding now with our approach, the constraint Hamiltonian is given for $(x,z,S)\in\mathbb{R}^{2}\times\mathbb{R}^{\star}_{+}$ by
${\cal H}^{\circ}_{\rm A}(x,z,S)$ $=\kappa x-\inf_{a\in\mathbb{R}_{+}^{d}}\big{\{}a\cdot\mathbf{1}_{d}z+c_{1}(a)\big{\}}-\dfrac{1}{2}\inf_{b\in\mathbb{R}_{+}^{d}\text{ s.t }\sigma(b)\sigma(b)^{\top}=S}\big{\{}c_{2}(b)\big{\}},$
and leads to the following optimiser
$a^{\circ}_{k}(z)=-\mu_{k}(z\wedge 0)\;\text{ and }\;b^{\circ}_{k}(S)=\frac{S}{\widebar{\sigma}\sqrt{\lambda_{k}}}.$
We can then compute the dynamics of $X$ and $Y^{y_{0},Z}$ under the optimal effort, and rewrite the principal’s problem as a more standard control problem, similar to (4.3). It is nevertheless already obvious that, to have equivalence between the two problems, we need to ensure that the quadratic variation of $X$ in both problems coincides, leading to the following relationship between $S\in\mathbb{R}_{+}^{\star}$ and $\gamma\in\mathbb{R}_{-}$:
$\sum_{k=1}^{d}\sigma_{k}^{2}b_{k}^{\circ}(\gamma)=\sum_{k=1}^{d}\sigma_{k}^{2}b_{k}^{\star}(\gamma)\quad\Leftrightarrow\quad S=\dfrac{\widebar{\sigma}}{\sqrt{-\gamma}}\quad\Leftrightarrow\quad\gamma=-\dfrac{\widebar{\sigma}^{2}}{S^{2}}.$
One can verify that this coincides with the correspondence given by Lemma 3.11. Indeed, we have
$\operatorname*{arg\,max}_{S>0}\bigg{\{}{\cal H}_{\rm A}^{\circ}(x,z,S)+\dfrac{1}{2}\gamma S\bigg{\}}$ $=\operatorname*{arg\,max}_{S>0}\bigg{\{}-\frac{\widebar{\sigma}^{2}}{S}+\gamma S\bigg{\}}=\dfrac{\widebar{\sigma}}{\sqrt{-\gamma}},\quad\gamma<0,$
[…]
5 Conclusion
In this paper, we develop a novel, simpler approach to solving principal–agent problems with volatility control. Specifically, while the existing approach, introduced by Cvitanić, Possamaï, and Touzi (2018) [9], relies on the theory of second-order BSDEs, the approach we propose allows us to solve this type of problem solely using BSDEs. More precisely, in [9], the main and challenging result is to prove the equivalence between the original principal’s problem, $V_{\rm P}$ in 1, and its reformulation as a more standard stochastic control problem, $\widetilde{V}_{\rm P}$ in 2. While it is clear from the restriction to contracts of the form (2.15) that $V_{\rm P}\geq\widetilde{V}_{\rm P}$, the converse inequality requires to prove that such restriction is without loss of generality, which necessitates the use of second-order BSDEs. Even though this approach indeed allows a thorough analysis of the principal–agent problem with volatility control defined by [9], and therefore of the applications arising from this framework, the mathematical complexity of this approach hampers the study of natural and relevant extensions, such as to multi-agent frameworks. Moreover, in our opinion, this mathematically challenging approach also prevents a wider dissemination of these models and their relevant applications in many fields, from operation research to economics.
In contrast, we introduce in this paper an alternative ‘first-best’ formulation in 3, in which the principal directly controls the density of the output’s quadratic variation. This problem is purely theoretical: it can be seen as a dual to 1, and is not necessarily meaningful in terms of application. However, compared to 1, 3 can be solved in a more straightforward way, following the so-called Sannikov’s trick, as introduced by Sannikov (2008) [29]. More precisely, one can prove using BSDEs that the restriction to contracts of the form (3.12) is without loss of generality, which implies that the principal’s problem in this ‘first-best’ reformulation is actually equivalent to a more standard stochastic control problem, namely $\widetilde{V}_{\rm P}^{\circ}$ defined in 4. It then remains to show that the value $V_{\rm P}^{\circ}$ of the ‘first-best’ reformulated problem, which naturally satisfies $V_{\rm P}^{\circ}\geq V_{\rm P}$, can be achieved. For this, it suffices to use the contract form (2.11) considered in [9], and thus show that 2 and 4 are actually equivalent. In other words, instead of choosing the sensitivity parameter $\Gamma$ in the contract form (2.11), one can directly optimise the density process $\Sigma$ of the output’s quadratic variation. Conversely, given a density process $\Sigma$, one can recover from the one-to-one correspondence in Lemma 3.11 the appropriate pay-for-performance parameter $\Gamma$ in the contract form (2.11). We finally obtain that the problems introduced throughout this paper are all equivalent, which is graphically summarised by Figure 1.
1234Lemma 2.5, or [9]Lemma 3.4Theorem 3.9By 2BSDEs in [9]Theorem 3.10Lemma 3.11
Figure 1: Equivalence of the problems
We nevertheless wish to mention that, at the end, and depending on the framework under consideration, 2 may still be easier to solve than 4. Indeed, as illustrated through Section 4, the agent’s problem 4 involves a constrained optimisation, making it potentially more complex to solve in practice. However, this does not undermine the effectiveness of our approach, since its main goal is to provide a theoretically simpler method to establish the optimality of the relevant contract forms. Ultimately, since all problems are equivalent, this approach offers the flexibility to choose the most practical problem to solve. We also wish to emphasise once more that the contract form (2.11) introduced in [9] is essential for proving the equivalence between the original problem and its ‘first-best’ reformulation. Nevertheless, the key message of our approach is that we do not need to rely on 2BSDEs to demonstrate the optimality of these contracts; their optimality is ensured by the fact that the ‘first-best’ value is achieved.
We believe that this alternative, simpler approach will provide a more straightforward pathway for extending principal–agent problems to more complex frameworks, involving for example multi-agent settings, mean-field interactions, or non-continuous output processes. For instance, solving principal–multi-agent problems with volatility control via the original approach in [9] requires the consideration of ‘multidimensional 2BSDEs’, as highlighted in [18], for which the theory is actually not yet developed. In contrast, following our proposed approach, one could rely on the more mature theory of multidimensional BSDEs to show the equivalence between the original problem and its ‘first-best’ reformulation, thereby ensuring optimality of appropriate revealing contracts without relying on multidimensional 2BSDEs. Furthermore, this simplified approach shall facilitate practical and more widespread implementation of principal–agent problems with volatility control in diverse fields, beyond mathematical finance. Indeed, such models are already used notably for energy-related applications, see [1, 11, 2], and may also prove useful in other domains.